A Complete Bibliography of ACM ForTec Forum and ACM Fortran Forum

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 March 2021 Version 2.50

Title word cross-reference

1000! [Ola01a].

007r1 [ANS95].

1 [NAG85]. 1.1 [Ano01c]. 10 [Rei90]. 11 [Ano85p, CS12b, Kha91]. 12 [Ano85p, Rei89d]. 16 [CS15b, Rei90]. 17 [Rei90]. 19 [CS16b]. 1978 [Com89]. 1982 [Ano82a]. 1984 [Ano85p]. 1985 [Ano85p, Ano85q, Ano85n, NAG85]. 1986 [Rei86b, Rei86a, Rei86c, Rei87a, Rei87f]. 1987 [Miy87, Rei87c]. 1989 [Com89, Mei89b]. 198X [Com89, Ano87d]. 1990 [Ano90i]. 1992 [Ano92r, Bro92]. 1993 [Ano93e, Hig94a, Mei92e]. 1995 [ANS95].

1998 [Ano97b].

2 [Cla83, Mei90e]. 20 [CS16c]. 2000 [Ano95l, Ano99a, Ano00e, Koe98b, Koe98a, Mei96b, Nag03, Rei00a, Rei00b, Rei02a, Rei02c, Rei03, Ano97d, Wag95]. 2001 [Ano00a, Hol00]. 2002 [Coh97, Coh99, Don02]. 2003 [CL10, Car10, CS07a, CS07b, CS08a, CS08b, CS09b, CS09c, CS10c, CS11a, CS11b, CS11c, CS12c, CS12a, CS13b, CS14, CS15a, CS16a, CS18b, Don06, LC11b, Mar03b, Mar08a, PCV12, Rei07]. 2004 [Sny04]. 2006 [Bad07a]. 2008 [Ano06a, CS09a, CS10b, CS13a, Don07, Mux09, Nag09, Rei08, Rei14, SMCH15]. 2009 [Mux09, Nag09]. 2015 [Rei17]. 2018 [CS18b, Rei18]. 21 [CS17a]. 22 [CS17b, CS18c]. 22.02.01.04 [ISO96b]. 24 [CS18a]. 25 [CS18b]. 26 [CS19a].

3 [Ano83b].

4.0 [Ano95o, Ano97k]. 4th [Iwa00].

5.0 [Ano97j]. 55 [Mei89b].

6 [Ano85q, CS10a]. 6.0 [Ano01c].

754 [Bie95]. 77 [Ano82c, Ano83a, Ano84e, Ano84j, Ano85e, Ano85m, Ano92q, Ano93d, Ano94o, Ano94n, Ano94l, DTF92, Mei83a, Mei84a, Mei84b, Met83a, MCL^{+}95, Pat93, Red91, Val83, Wag83b, Wag83a, Ano85h, Ano85i].

8 [Rei89e, Rei90]. 82/18 [Ano83b]. 82nd [Ano82a]. 83 [Ano83b]. ’84 [Ano85j]. 88 [Mei89b, Mei89a, Mei89d]. 8X [AR87, Ano83a, Ano83b, Ano84j, Ber85b, Com89, Dol88, Ful87, Lah87, Mei87c, Mei88a, Met87, Miy87, Rag82, Rei89a, Rei89b, Smi85, Ano85o].

90 [Ano91g, Ano92p, Ano92q, Ano92r, Ano92s, Ano93d, Ano94o, Ano94n, Ano94m, Ano94q, Ano94v, Ano94u, Ano95g, Ano95k, Ano95o, Ano95p, Ano96h, Ano96i, Ano96l, Ano96j, Ano97a, Ano97h, Ano97i, Ano97n, Ano97j, BGvE^{+}97, Ble96, DNS97, DNS99, DN99, DN00a, Dec08, Dup94, EEV^{+}96, Kai03b, Lau92, Mar01, Mar06, Mar92, Mei90a, Mei90c, Mei90b, Mei91a, Mei91b, Mei92a, Mei92c, Mei92b, Mei92d, Mei93, Mei97c, Mei97b, Met99c, Met99d, Met99e, Met00c, Met00a, Met00b, Met01c, Met01d, Met01b, Met02b, Met02c, Mit99, MCL^{+}95, OAB^{+}95, Pre95, Pri99, Red91, Red94, Rei93a, Sch93, Sny97, SNB^{+}98, Tay97, Wag93, Ano91g, Ano95b, Cis94]. 90/95 [Mar03b, Ano96h, Mar06, Tay97]. 90/95/HPF [Met99c, Met99d, Met99e, Met00c, Met00a, Met00b, Met01c, Met01d, Met01b, Met02b, Met02c]. 95 [Ano93f, Ano94v, Ano94u, Ano94x, Ano96h, Ano97b, Ano97k, Ano98b, Bad06, Bad07a, BGvE^{+}97, Ble09, CL10, Car10, Cle99, Ded95, Don02, FRM^{+}97, Hat03, Mar06, Mei94c, Met96, Met06, MCL^{+}95, OFKJ10, RW99b, Sch01, Sny00, Tay97, Wag95, ANS95]. 95-007r1 [ANS95]. 95/2002 [Don02]. 95/2003 [Sny04, CL10, Car10]. ’98 [Ano98d, Ano97b].

abridgement [Miy87]. Absoft [Goe08]. Abstracting [Met20]. Abstracts [CRC^{+}98]. Academic [Ano95p, BGvE^{+}97]. Acceptance [OAB^{+}95]. access [Bie95]. ACM [Ano90b, Ano90a, Ano91a, Ano94a, Ano95a, EEV^{+}96, Mei84b, Mei96c, Met99h]. ACM-Published [Met99h]. Actions [Ano82a]. Activities [Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano91j, Ano91k, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Ano97d, Clo99]. Activity [Ano00e]. Actual [CBFM^{+}98]. Adams [Ano85p, Ano91g]. Additional [Lon16]. Advance [EEV^{+}96]. Advanced [Ano98b, OAB^{+}95]. Akin [Mar03b]. Albuquerque [Rei87d]. Algebra [OAB^{+}95]. Algorithm [Hat03]. Algorithms [Ano96h, Ano00c]. Aliasing [CBFM^{+}98, Gil96]. Alive [MCL^{+}95]. Alliance [Ano97b]. Allocatable [BGvE^{+}97, FRM^{+}97, ISO96b, Mai96, Mei95b, OAB^{+}95]. almost [Sch93]. alternate [Ful87]. Alto [Rei89d]. Amazing [Mei92a]. Amendment [Mei84b]. amendments [Wag93]. American [Ano87d, Com89]. Analysis [Ano95m, Ano01b, RW99a]. Annals [AFI84, Ano82c]. Announce [Ano97b]. Announced [Mei89a]. Announcement [ACM94, AWP82, Ano98a, Ano99a, Mei92a]. Announcements [Ano82d, Ano83a, Ano83g, Ano83h, Ano84a, Ano90b, Ano90a, Ano91a, Ano92a, Ano92b, Ano95a, Ano96l, Ano97a, Ano97b, Ano98b, Ano99d, Ano99e, Ano00b, Ano01a, Ano01b, Ano01c, Ano02a, Ano02b, Ano03a]. Announces [Ano87d]. Annual [Iwa00]. ANSI [AWP82, Ano82a, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Mei92a, Miy87]. APLs [She00]. appendix [Ded95]. Applications [BGvE^{+}97, Kuh98, Ano85j]. Applied [Ano01c]. Approach [Don01a]. Approval [Mei92a]. Approved [Ano97k]. April [Ano90i, Ano96i, Rei86a]. Archaic [Nag05, Nag06]. archive [Wag03]. Argonne [Ano85j]. Argument [CBFM^{+}98, Mei95b, Rag82]. Arguments [ISO96b, MCL^{+}95, Ano97d, Wag95]. Arithmetic [Ano98b, Del00a, DTSB93, Mei95b, Sch01]. Arjen [GD06]. Array [BGvE^{+}97, ISO96b, Ngu03, Cle09, Mar08b, NR98]. array-valued [Mar08b]. Arrays [BGvE^{+}97, FRM^{+}97, Hen90, ISO96b, Mei95b, Rag82, Don07, Mai96, NR05, OAB^{+}95, Red94, Rol95]. article [GD06]. Articles [Ano84g, Ano90f, Ano90e, Ano91f, Ano91g, Ano92h, Met99h, MCL^{+}95]. Association [BGvE^{+}97, Hig94d, Rag82]. Attribute [BGvE^{+}97, CBFM^{+}98, Wag95]. Aug [Ano83b, Rei89c]. August [Ano00e, Rei00b, Ano90i, Ano96l, Bad07a, Rei87a, Rei87b]. Austria [Rei90]. Authors [Met99g, Met99h]. AutoFortran [ALN18]. Automated [Aha85a]. Availability [FGMS91]. Available [MCL^{+}95]. Avoiding [Mar03a]. avoids [Mei89e]. Award [Ano94b, Sof99]. Awards [Ano90b, Ano90a, Ano91a]. aware [CS05a]. away [Met20]. baby [Mei90d]. Back [Bee02]. Backus [Ano94b]. Bad [BGvE^{+}97]. ballot [Rei03]. Based [Hat03, Jon95, She00]. basic [Nag05]. Be [Jon95, Koe98a, Mei91b, MCL^{+}95]. becomes [Mei84a]. Benchmarks [Ano95n, Pol13]. beta [Ano01c]. Betty [BGvE^{+}97]. Between [Ano95a, Ano96j, Dec08, Red91]. big [Wag85b, Wag85a]. Binding [Ano89k, Bad07b]. Bindings [Ano98b, Mit99]. Blanks [BGvE^{+}97]. Bonn [Ano97b]. Book [Ano85i, Ano90f, Ano91g, Ano93c, Ano95b, Cis94, Met20]. Books [Ano90f, Ano90e, Ano91f, Ano91g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano94h, Ano94i, Ano94k, Ano95k, Ano97h, Ano97i, Met99e, Met02c, Met01c]. Boston [Rei89c]. Both [Ano83a]. Bound [FRM^{+}97]. Brainerd [Ano91g]. Brian [BGvE^{+}97]. British [Ano83a]. builder [CS07c, CS15c]. Business [Ano85j].

C [ALN18, Ano96f, Ano97b, Bad13, Dec08, Don06, EHH09, EEV^{+}96, FGMS90, FGMS91, Gil96, OAB^{+}95, She99, She00, Wag95]. CA [Rei89d]. Call [Ano83a, Ano99a, Mei98a]. callable [She99]. Cambridge [Mar03b]. can [Bad06, Ber85b]. capabilities [Pol12]. capability [CL10, CS19b, Chi20]. Captured [BGvE^{+}97, BRZ^{+}97, CBFM^{+}98, EEV^{+}96, FRM^{+}97, MCL^{+}95, dPGS^{+}96]. Cascade [SG02, Chm01a, Chm01b]. Case [FRM^{+}97]. Catalog [Ano85h]. CATO [Hat03]. CD [Rei03]. Celebrated [AFI84]. Certified [Ano84b]. Chairman [Wag82a]. challenges [Hey98]. Change [Mei94a]. Changes [Ano84j, AW89]. Changing [Mei82a]. characterization [Kai03b]. China [Ano83a]. Classics [Ano95k, Ano96h, Ano97h, Ano97i]. Close [Mei96b]. CMPLX [Vow01]. Co [NR05, Don07, Rei90, NR98]. Co-Array [NR98]. Co-arrays [NR05, Don07]. Coarray [FR15, Sie16]. Coarrays [Rei10, SMCH15]. Coast [Mei98b]. CoCo [Eps99, Ano01c, Nag01]. coco.f90 [Ano01b]. Code [Ano95m, Mar09, Kha91, TM12]. Coding [Met83a, Caf84]. Cohen [Sny04]. collection [OFKJ10]. Combining [Mar02]. Come [Ano87d]. Comment [Ano87d]. Comments [Ano84j, GD06, Rei89a, Rei89b]. Committee [AWP82, Ano82a, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Rei00b, Rei02b, Wag95, ANS95, Rei01]. Committees [Ano94t, Ano94u, Ano95i, Ano95h, Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Clo99, Ano97d, Ano93f, Ano94v]. common [Ano85j]. Communicating [Don01a, Don01b]. comp [EEV^{+}96]. comp-fortran-90 [EEV^{+}96]. comp.lang.fortran [Ano09]. Company [Ano01c]. Compaq [Ano01b, Ano01c]. Comparing [App01]. Comparison [Ano96j, Red91]. Comparisons [App14, App15, App16, Pol13, Pol12]. ’compilation [Chm01b, Mei96b, OAB^{+}95, Sch99, SG02, Chm01a, Mei96a]. Compiler [Ano95p, Ano98b, Ano00c, App14, App15, App16, CS06a, CS07a, CS07b, CS08a, CS08b, CS09b, CS09c, CS09a, CS10a, CS10b, CS10c, CS11a, CS11b, CS11c, CS12b, CS12c, CS12a, CS13b, CS13c, CS13a, CS14, CS15b, CS15a, CS16a, CS16b, CS16c, CS17a, CS17b, CS18b, CS18c, CS18a, CS19a, CS20, Mei91a, Val83, Wag83b, Ano01c, Pol13, Pol12]. Compilers [Ano84b, Ano95n, Ano95m, Mei92c, Mei92b, Mei92d, App01, CS06b, CS19b, Chi20, Kai03b, Met99c, Met00a, Met01b, Ver90]. Compiling [Bus17]. Complete [Mei91a]. Components [Don02, ISO96b, Mai96]. Composite [Ano85n, Ano85o]. Computational [Ano00a, Ano01c]. Computer [Ano83a, Ano85i, Ano85j, Ano90b, Hat03, Gui87, Ano84g]. Computers [Wag83b]. Computing [Ano98d, Ano01d, Mar04b, AFI84]. concepts [DNS97]. Concerning [Mei83b, Mei87c, WRV^{+}92, Tay97]. Conditional [Mei96b, OAB^{+}95, Mei96a]. Conference [Ano82b, Ano83a, Ano94a, Ano00a, Mei89a]. Conferences [Ano83c, Ano83d, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano87a, Ano87b, Ano87c, Ano88a, Ano89a, Ano89b, Ano89c, Ano90c, Ano90d, Ano91b, Ano91c, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano93b, Ano94e, Ano94f, Ano94g, Ano95c, Ano95d, Ano95e, Ano95f, Ano96a, Ano96b, Ano97c, Ano97h, Ano98c]. Confused [MCL^{+}95]. consecutive [Gor08]. Conservative [Run95]. Considerations [Hen90]. Consortium [Ano94e, Ano94w]. Constitutional [Mei84b]. Construct [Ano85n, Ano85o]. Constructing [Sny00]. Constructive [Sny01]. Constructor [BGvE^{+}97]. Consultancy [Met03b]. contd [Bee01]. Content [Rei00a, Ano97d]. Contents [NAG85]. Contiguous [Hen90]. Continue [Cla83]. Contributing [Met99g]. Convenor [Rei99, Rei00b, Rei02b]. Conventions [Ano85i, Met83a, Caf84]. Conversion [Ano97j, Aha85a, Met91]. converter [FGMS90, FGMS91]. Converters [Ano95q]. Copies [Ano92r, Ano93f, Ano94p, Ano94u]. Corp. [Ano95o, Ano97k]. Corporation [Ano86f]. Correction [Bad07a, DN00b]. Correspondence [Ada82, Ano93a, Ano94c, BPP^{+}82, BFP83, BJ90, Ber85a, Bro92, Clo94, DB85, Del93, DTSB93, Ell93, HBM87, Hen90, Her94, Hyb87, IMH85, KH85, Lau92, MLW84, Mei90a, Met91, Met94, MF92, Mur89, OCMB94, PBF86, Rit87, SHK83, SWM^{+}89, SH89, SBDB87, TV83, TSSL88, WM83, WRV^{+}92]. Corresponds [CBFM^{+}98]. Count [Mei92a]. counting [Car10, LC11b]. Courses [Ano97h, Met01d, Met03b, Met00c]. Coverage [Ano01b]. Cray [Ano94d, CS19b, Chi20, Mar92]. creating [Cle09, Kai03a]. Cross [Sin99]. Cross-platform [Sin99]. Crunch [Vau00]. Current [Ano91j, Ano91k, Mei90a].

Damian [Mar14a]. dark [Mee90]. Data [DTF92, Wal00, Dec08, Gil01, Gor14a, Gor14b, SM87, BGvE^{+}97]. data-type [SM87]. date [Wag93]. David [Nag09]. Day [Mei92d]. Days [Mei87a, Ada86]. deadlock [Mei89e]. Debunking [Ded99]. DEC [EEV^{+}96]. DECUS [EEV^{+}96]. deep [Cur18]. default [Whi00]. Defense [Mei92c]. defined [Mar14b]. Degenerate [Red94]. Delft [Mux13]. Department [Mei91a, Mei92c]. Deplore [EEV^{+}96]. derived [ISO96b, Mar03a, Ste03]. Describe [Mei95b]. Descriptors [Rol95]. Design [Ano85i, GD06, Mar06, Mar08a, Ful87, Mar14a]. Desired [Ano83b]. Developer [Ano97k]. developing [CS06a]. Development [Ano95a, Goe08, MDB93, CS05a, Lut95, Mei98b, Sin99]. Developments [Mei82b]. diagnostic [CS19b, Chi20, Pol12]. dialect [BCS92]. DIGITAL [Ano97j, Ano95o, Ano97k]. Dinosaur [Mei94b]. Dinosaurs [OCMB94]. Disciplined [Ano96k]. discussion [Dol88]. discussions [Ano99c]. Distribution [Mei83b, Mei90d]. Do [Ano96h, Mei88b, Ber85b, Red93, Rei93b, Rei93a, Ano85n, Ano85o, BGvE^{+}97]. Do-it-yourself [Ano96h]. Document [Miy87]. Documentation [Ano83a]. Documents [Ano97h, Ano98b]. Donation [Ano01b]. Don’t [BGvE^{+}97]. Double [Mei90d]. Downloadable [Ano98b, MCL^{+}95]. Draft [Ano95n, Bro92, EEV^{+}96, Mei94c, Mei96d, MCL^{+}95, ANS95, AR87, AW89, Com89, Ded95, Mei87c, Mei90d, Ano87d]. Dummies [CBFM^{+}98]. Dummy [ISO96b, Mei95b, Ano97d, Wag95]. Dvorak [Ano85p]. Dynamic [Bad06, Bad07a].

ECMA [Ano83b]. ECMA/TC8/83/3 [Ano83b]. Ed [Mar03b]. Editor [Mei91a, Mei93, Mei82a, Mei83a, Mei83b, Mei83c, Mei84a, Mei84b, Mei85a, Mei85b, Mei85c, Mei90c, Mei90d, Mei91b, Mei92a, Mei92c, Mei92b, Mei92d, Mei92e, Mei94b, Mei94a, Mei94c, Mei95b, Mei95a, Mei96c, Mei98a, Wag82b, Wag83b]. Editorial [Mei89b, Mei89a, Mei96b, Mei98a, Mei99, Met99a, Met99b, Met01a, Met02a, Met03a, Ada86, Mei85d, Mei86, Mei87c, Mei87a, Mei87b, Mei88b]. editors [CS05a]. Educational [Ano01b]. Edward [Rei89a]. Effective [Ano85h, Ano85i]. Effects [FRM^{+}97]. Efficiency [BGvE^{+}97, Hen90]. Efficient [MCL^{+}95, Kha91]. Electronic [ACM94, Ano92s]. Elemental [Wag95]. ELF [Ano96j, Ano95p, Ano96h, MCL^{+}95, Run95]. ELF90 [Ano96l, Ano97h, Ano96k]. Ellis [Ano95b]. Email [MCL^{+}95, OAB^{+}95, FRM^{+}97]. emerging [Met87]. Enable [OAB^{+}95]. encapsulation [DTF92]. encounter [Met92]. end [Bus10]. Engineering [Ano97k, Ano98b, Ano01c]. English [Met99e]. Enhanced [Ano98b, ISO96b]. enough [Whi00]. Enumerated [Gil01]. Environment [Ano85m, Ano85j, Goe08]. environments [CS05a]. Equipment [Ano95o, Ano97k]. Equivalent [BGvE^{+}97]. Errata [Wag93]. Error [Eps96, Eps97, PCV12]. Esprit [Ano85j]. Essential [Ano96j, MCL^{+}95]. Estes [Rei90]. etc [Met00c, Met01d, Met02b]. Europe [Ano99a, Hol00]. European [Ano85j, Wag82b]. evolution [Sch91, Van84]. Example [Sie16]. examples [MCC96]. Exception [Mar13, Rei95, Sny15, Sny16, ISO96a]. exceptions [Bie95]. Excerpts [Ell93]. Execution [Ano97h, Pol13]. EXIT [Rei93b]. Experience [Ano97j]. Experimenting [Mar19]. Expert [DTSB93]. Explained [Sny04, Ano91g]. Expressing [DNS97]. Expressions [BGvE^{+}97, DTSB93]. Extended [Bad13]. extending [Don04]. Extension [Ano97d]. extensions [Lev94, The91, Rei92]. External [Mei95b]. Extrinsic [Hig94c].

F [Cis94, Ano96h, Ano96j, Ano96k, Ano97h, Bra96, Eps96, Eps97, FRM^{+}97, Met98]. F2000 [Mei96b]. F2005 [FRM^{+}97]. f2c [Ano95q, FGMS91]. F2X [Don01b, Don01a]. F77 [Ano95m, Ano95q, Mei95b]. F90 [Ano95a, Ano95n, Ano95q, Mei95b, Pag96, Sch99, Met92]. f90SQL [Ano01b]. F95 [Ano95n, Don01b, Ano01c]. Faces [Mei91a]. facilitating [ALN18]. Facilities [Del00a, ISO96b]. Facility [Ano95l, Bur87, Mei96b]. Failure [Ano83a]. Fall [EEV^{+}96, Mei89b]. Fast [Mei86]. FCAT [Ano01b]. Feature [Ano82c]. Features [Ano96j, Lon16, Mei90a, Rei02c, Ded95, Mar19, Rei06, Rei07, Rei08, Rei14, Rei17, Rei18]. Feb [Rei89d]. February [Ano85p, NAG85, Rei87c, Rei88]. File [Ano97n, Met99e, Met01d, Met02b, Met02c, Met03b, Met99c, Met99d, Met00c, Met00a, Met00b, Met01c, Met01b]. files [Mar04a, Mar05]. Final [Ano91g, EEV^{+}96, Ano97d]. Find [MCL^{+}95]. Finishing [Mei90d]. Finland [Ano00e]. First [Ano99a, Mei91a, Met92]. flaw [Kha91]. FLIB [Ano95m]. Floating [OAB^{+}95, ISO96a]. floating-point [ISO96a]. Flowchart [Ano95m]. Follow [Ano85e]. Follow-up [Ano85e]. Font [Mei94c]. foresight [Mee90]. ForeSys [Ano96l]. Form [Ano92s, BGvE^{+}97, Mei93, Mei90b, Met91]. Formal [Ano95a, MCL^{+}95]. Forpedo [McC05]. Fort [Rei87g]. ForTec [Mei83c, Mei84a, Mei84b, Wag82a]. Forth [Ano89l]. Forthcoming [Ano83c, Ano83d, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano87a, Ano87b, Ano87c, Ano88a, Ano89a, Ano89b, Ano89c, Ano90c, Ano90d, Ano91b, Ano91c, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano93b, Ano94e, Ano94f, Ano94g, Ano95c, Ano95d, Ano95e, Ano95f, Ano96a, Ano96b, Ano97c, Ano98c, Mei92c, Mei92b]. Fortran [AWP82, Ano82a, Ano83a, Ano83b, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano84h, Ano85h, Ano85i, Ano85j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano91g, Ano94v, Ano95b, Ano95o, Ano98a, Ano00e, Ano01b, Ano01c, Bad07a, Bur09, Com89, CS15c, Cis94, Goe08, Mar03b, Mei84b, Met20, MCL^{+}95, Rei00b, Rei02b, Sny04, EEV^{+}96, AFI84, AR87, AW89, Aha85a, ALN18, Ano82c, Ano83a, Ano83b, Ano84b, Ano84e, Ano84g, Ano84j, Ano85e, Ano85i, Ano85n, Ano85o, Ano85m, Ano86f, Ano87d, Ano89k, Ano91d, Ano91g, Ano92k, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano92v, Ano92s, Ano93e, Ano93d, Ano93f, Ano94o, Ano94n, Ano94m, Ano94l, Ano94p, Ano94q]. Fortran [Ano94s, Ano94v, Ano94u, Ano94x, Ano94y, Ano95a, Ano95g, Ano95i, Ano95h, Ano95k, Ano95l, Ano95n, Ano95o, Ano95p, Ano96c, Ano96d, Ano96e, Ano96f, Ano96h, Ano96i, Ano96l, Ano96j, Ano96k, Ano97a, Ano97b, Ano97e, Ano97f, Ano97h, Ano97i, Ano97n, Ano97l, Ano97m, Ano97j, Ano97k, Ano98b, Ano98d, Ano98e, Ano98f, Ano98g, Ano98i, Ano99b, Ano99c, Ano00c, Ano00d, Ano00e, Ano01c, Ano01e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Ano06a, App01, App14, App15, App16, Arm12, Bad06, Bad07b, BCS92, BGvE^{+}97, Bee01, Bee02, BRZ^{+}97, Ber85b, Bie95, Ble96, Ble09, Bro92, Bus17, Bus10, Com89, Caf84, CL10, Car10, CBFM^{+}98, CS05a, CS06a, CS06b, CS07a, CS07b, CS07c, CS08a, CS08b, CS09b, CS09c, CS09a, CS10a]. Fortran [CS10b, CS10c, CS11a, CS11b, CS11c, CS12b, CS12c, CS12a, CS13b, CS13c, CS13a, CS14, CS15b, CS15a, CS16a, CS16b, CS16c, CS17a, CS17b, CS18b, CS18c, CS18a, CS19a, CS19b, CS20, Chi20, Cla82, Cla83, Cle99, Clo94, Clo99, Coh97, CRC^{+}98, Coh99, Cur18, DTF92, DNS97, DNS99, DN99, DN00a, Dec08, Ded95, Ded99, Del93, Dol88, Don02, Don04, Don06, Don07, Dup94, EHH09, Ell93, EEV^{+}96, Ell98, FR15, FGMS90, FGMS91, FRM^{+}97, Ful87, Gil96, GFT^{+}99, Gor08, Gor14a, Gor14b, Gre98, Gui87, Hig94a, Hig94b, Hig94c, Hig94d, Hat03, Hen18, Hey98, Hir96, Hol95, Hou05, Iwa00, Jon95, Kai03b, Kha91, Koe92, Koe98b, Koe98a, Kuh98, Lah87, Lau92, Lev94, LC11a, LC11b, Lon16, Mar01, Mar02, Mar04b, Mar06, Mar08a, Mar13, Mar92, Mar03c, McC05]. Fortran [Mee90, Mei82b, Mei83a, Mei83b, Mei84a, Mei84b, Mei85d, Mei87c, Mei88a, Mei89b, Mei89a, Mei89d, Mei90a, Mei90c, Mei90d, Mei90b, Mei90e, Mei91a, Mei91b, Mei92a, Mei92c, Mei92b, Mei92d, Mei92e, Mei93, Mei94b, Mei94a, Mei94c, Mei95b, Mei96c, Mei96b, Mei96a, MCC96, Mei97c, Mei97a, Mei97b, Mei98a, Mei98b, Met83a, Met83b, Met84, Met87, MR91, Met95, Met96, Met98, Met99c, Met99d, Met99e, Met00d, Met00c, Met00a, Met00b, Met01e, Met01c, Met01d, Met01b, Met02b, Met02c, Met02d, Met03b, Met06, Met19, Mit99, Miy87, MCL^{+}95, NAG85, Nag02, Nag03, Ngu03, NR98, NR05, OAB^{+}95, OFKJ10, The91, Pag05, Pat93, Pol13, PCV12, Pre95, Pri99, Rag82, Red91, Red94, Rei89a, Rei89b, Rei92, Rei93a, Rei95, RW99a, RW99b, Rei00a, Rei01, Rei02a, Rei02c, Rei03, Rei06]. Fortran [Rei07, Rei08, Rei10, Rei14, Rei17, Rei18, Ros95, Run95, Sch91, Sch92, Sch93, Sch01, SWM01, She99, She00, SMCH15, Sie16, SVM^{+}93, Sin99, Smi85, Sny97, Sny00, Sny15, Sny16, Pol12, Ano97d, SNB^{+}98, Tay97, TLBC00, TM12, Val83, Van84, Vau00, Ver90, Vow99, Wag83b, Wag83a, Wag84, Wag85b, Wag85a, Wag93, Wag95, Wag97a, Wag97b, Wag97c, Wag03, Wal00, WRV^{+}92, ANS95, GD06]. Fortran-Hating [Del93]. Fortran-Oriented [Ano96h, Ano96i]. FortranPlus [Ano01c]. Fortrans [Aha85b]. FORTVER [Ano86f]. Forum [Bro92, Rei02a, Ano01e, Bee02, Mei84a, Mei92c, Mei92b, Mei92e, Mei94a, Mei96c, Mei98a, Met00d, Met01e, MCL^{+}95, Wag82a, WRV^{+}92]. FPP [FS01, Ano96k]. Framework [MDB93, Cur18, LC11a]. Francais [Ano96h]. France [Ano82b, Ano99c, Met99f]. Free [Ano96l, Ano01c, Mei93, MCL^{+}95, Vau00]. freeware [Ano95m]. French [Ano98a]. F^{T M} [Ano96j]. FTN90 [Ano95o]. ftp [EEV^{+}96]. Fujitsu [Ano97b, Ano98b, Ano01c]. Full [FS01]. Function [BGvE^{+}97, ISO96b, Vow01]. functionality [SM87]. Functions [FRM^{+}97, Jon95, OAB^{+}95, Mar08b, She99]. Future [Ano83b, Ano90i, Ano91j, Ano91k, Ano92r, Ano92t, Ano92v, Ano92u, Ano92s, Ano93f, Ano94r, Ano94u, Ano95i, Ano95h, Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Clo99, Ell98, GFT^{+}99, Mei82b, Mei87a, Ano97d, Ada86, Aha85b, Ano99c, Cla82, Cla83, Met84, Sch91, Ano85i]. Futures [Ano97b, Ano98d, Ell98, CRC^{+}98, Gre98, Koe98b, Kuh98, Met98].

G [AWP82]. g77 [Ano95q]. Gateway [Del93]. Geezers [Hen18]. General [ACM94]. Generate [Mei93]. generation [Kha91]. Generations [Ano85i]. Generators [Ano95m, BRZ^{+}97]. Generic [BGvE^{+}97, Jon95, Mar01, McC05, Ano97d, Vow01, Ble09, Mar19, MY01]. Genericity [Don04]. Germany [WCH^{+}03]. getting [Wag85b, Wag85a]. gfortran [CS19b, Chi20]. giants [Mei97a]. Global [FRM^{+}97]. Glorious [Mei87a, Ada86].

GNU [Ano95q, Bad07b, Vau00]. Go [Mei88b]. going [Mei90d]. Goldberg [Ano91g]. Good [Mei87a, Ada86, Whi00]. Gotcha [Cle99, DN00b]. Gotchas [DNS99, DN99, DN00a]. Grammar [Ros95]. Grammars [MCL^{+}95]. Grin [Her94]. Ground [Mei99]. Group [Ano89l, Ano98b, Ano98a, FRM^{+}97, Iwa00, Mei89b, Mei92c, MCL^{+}95, Wag95, Ano89k, Mei98b]. Groups [Ano92r, Ano92s, Ano93f, Ano94s, Ano94u, Ano95i, EEV^{+}96]. Guide [Ano85j, Ano91g]. Guidelines [ACM94, Ano83a].

h2m [ALN18]. h2m-AutoFortran [ALN18]. Handling [Sny15, Sny16, Gil96, ISO96a, Mar13, Mar17, PCV12, Rei95]. Hardware [Ano85f]. Hating [Del93]. Heavens [Bee01, TLBC00]. HECToR [Bus10]. help [Bad06]. Here [Mei88b, Mei90d]. High [Ano98d, Bus10, Ano92k, Ano92v, Ano93e, Ano93f, Ano94v, BGvE^{+}97, Iwa00, Koe92, MCL^{+}95]. History [AFI84, Ano82c, Ano90b, Ell93, Mei98a, Met02d, Met19, Met20]. Ho-Hum [Mei92a]. Holberton [BGvE^{+}97]. Home [EEV^{+}96]. Honors [BGvE^{+}97]. Hot [Ano96h]. HPCN [Ano99a, Hol00]. HPF [Ano96i, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, EEV^{+}96, Met99c, Met99d, Met99e, Met00c, Met00a, Met00b, Met01c, Met01d, Met01b, Met02b, Met02c]. HPFF [Ano92r, Mei94a]. HUG2000 [Iwa00]. Hum [Mei92a]. HyperLiterature [Ano95a].

IBM [Ano85m, FR15]. Idaho [AWP82]. Idea [BGvE^{+}97]. IEC [Ano00e]. IEEE [Ano85j, Ano89k, Bie95, Del00a]. If [Del00b]. II [Coh99, Don01b, Hig94b, Hig94c, Hig94d]. III [Hig94a]. Image [Mei94b]. Implementation [Ano92r, Ano97b, Ble96, Hat03, Mei89a, Mei91b, Wal00, Ano96k, Ble09, Car10, LC11b, MY01, Nag01]. Implementations [Mei90a]. Implied [BGvE^{+}97]. Important [Mei89b, Gre98]. IMSL [Ano86f]. INCLUDE [Sny01]. Increasingly [Mei89b]. Indenting [Met02e]. Information [Ano82e, Ano87d, Ano90i, Ano90j, Ano91j, Ano91k, Ano92r, Ano92t, Ano92v, Ano92u, Ano92s, Ano93f, Ano94o, Ano94p, Ano94r, Ano94q, Ano94s, Ano94v, Ano94t, Ano94u, Ano94x, Ano94w, Ano94y, Ano95n, Ano95r, Ano96i, Ano96l, Ano97n, Ano97l, Ano97m, Ano98b, Ano98i, Com89, Mei94c, Met99e, Met01d, Met02b, Met02c, Met03b, Met95, Met99c, Met99d, Met00c, Met00a, Met00b, Met01c, Met01b]. informative [Ver90]. Infotron [Ano86f]. Initialized [Mei95b]. Input [Ano92a, Ano92b, Ano93a, Ano94c, Mei91a, Mei92c, Mar17, Mei89c, Pag05]. input/output [Pag05]. Institute [Ano86f]. INTEGERS [Wag83a]. Integrated [Goe08, MDB93]. Intel [Ano97k, Ano06c, CS19b, Chi20]. interchange [Ano09]. Interest [FRM^{+}97]. International [Ano83a, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano90i, Ano90j, Ano91j, Ano91k, Ano92t, Ano92v, Ano92u, Ano99b, Ano00d, Ano00e, Ano00a, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Clo99, Rei01, Rei02b, Wag95, Ano82b, Rei00b]. Interoperability [Don06, OAB^{+}95, Wag95, ALN18]. interoperation [Bad13]. Interpretations [Ano92r, Ano93f, Ano94q, Ano94v, Ano94u, Mei83a, Wag93]. Interval [Wal00]. Intrinsic [BRZ^{+}97, Hig94b, Gor08, Gor14a, Gor14b]. Introducing [Arm12, Wag82a]. Introduction [CS07c, CS15c, Mei96d, CS09d]. isn’t [Whi00]. ISO [Rei89c, Ano99c, Ano00e, Mei89b]. ISO/IEC [Ano00e]. ISO/WG5 [Rei89c, Ano99c]. Ispra [Rei90]. Issue [Ano82c, Ano93e, Hig94b, Hig94c, Hig94d, Mei92e, Mei96d, The91, Wag93, ANS95, Bad07a]. Issues [Ano84g, Bee02, Nag05, TM12]. Italy [Rei90]. Iterative [Don01b].

J3 [Ano06b, Nag09]. J3/WG5 [Nag09]. Jackson [Rei89c]. James [Cis94]. Jan [Rei90]. January [Rei86b, Rei86c]. Japan [Rei00b, Iwa00]. Jeanne [Ano85p]. Jefferson [Rei89e]. Jim [Mar14a]. John [Sny04, Ano85p, Ano85n, Ano94b, Mei90d, Rei86b, Rei86a, Rei86c, Rei87a, Rei87f, Rei87c, Rei87d, Rei87e, Rei87b, Rei88, Rei89c, Rei89d, Rei89e, Rei90]. Joint [Nag09, Mux13]. Journal [Ano83a, Ano85i, Ano01c, MCL^{+}95]. JTC1 [Ano00e, ISO96b]. JTC1/SC22/WG5 [Ano00e]. Jul [Rei90]. July [Rei90, Ano85n]. June [Bro92, Hol00, Miy87, Mux13]. Jurassic [Mei94c]. just [Eps99].

Kerrigan [Cis94]. Keynote [Ell98]. kill [Aha85b]. KIND [Whi00]. kinds [Nag06].

Lahey [MCL^{+}95, Ano95b, Ano95o, Ano96j, Ano97b, Ano98b, Ano01c]. Lahey/Fujitsu [Ano01c]. Language [Ano83b, Ano87d, Ano89l, Ano93e, Ano93f, Ano94v, Ano96j, Com89, Hig94a, Hig94b, Hig94c, Hig94d, Her94, Mei82b, Met20, CL10, EHH09, Koe98a, Lev94, Mei98b, Met98, Wag84, Ano84g]. Languages [Ano82b, Ano96h, Lut95, Mar02]. large [Kai03a]. Latest [Ano85o, MCL^{+}95]. Lauderdale [Rei87g]. leaks [Mar03a, Ste03]. Learn [Ano83a]. learning [Cur18]. Lecture [Ano94k]. legacy [LP00, TM12]. legitimacy [Met03a]. Len [Mos90]. Length [Bro92, Sch92]. Let [Aha85b]. Letter [Ano85p, Ano99c, Met99f, Wag82a]. Leveraging [FR15]. Lexers [Mei93]. LF90 [Ano95o]. LF95 [Ano01c]. Libraries [BGvE^{+}97, Ano95n, She00]. Library [Hig94b, Gre98, Pag96, Bad07b]. light [Mee90]. Likely [Mei89b]. Limited [Ano86f]. Line [Ano95o, Ano97h, MCL^{+}95]. Linear [Don01b, OAB^{+}95, Mei97b]. Linked [BGvE^{+}97, Ble09, Mei97c]. Linker [MCL^{+}95]. Linking [MCL^{+}95]. Links [Ano97h]. Linux [Ano01c, Ano95o, Ano96i, SNB^{+}98]. List [EEV^{+}96, Ble09, MY01, Mei97c]. Lists [BGvE^{+}97]. Lite [Ano95a]. Literature [Ano82c, Ano83e, Ano83f, Ano84f]. Little [Mei97a]. Liverpool [Rei87b]. Living [MCL^{+}95]. Local [CBFM^{+}98, Mei95b]. London [Ano97b, MCL^{+}95]. look [Mar04a, Mar05]. Loops [BGvE^{+}97]. Loren [Sof99]. Lower [FRM^{+}97]. LPM [Ano83b].

MA [Rei89c]. Machine [Ano91g, Met20]. Macintosh [Ano95m, Ano96i]. Magazine [Ano84g, Ano85j]. Mailbase [EEV^{+}96]. Maintaining [SVM^{+}93, Sny00]. Maintenance [SWM01]. major [Rei06]. MAKE [MCL^{+}95, Sny00]. Making [Del00a]. Malcolm [Sny04]. Mallory [Ano06b]. man [Sny97]. Management [Ano83a]. managing [Mei97c]. March [Mei92e, Rei87d]. Market [OAB^{+}95]. Markus [GD06]. Math [OAB^{+}95]. Mathematically [BGvE^{+}97]. Matrices [MCL^{+}95]. May [Ano82a, Ano82b, Ano85q, Ano90i, Hig94a, Hig94b, Hig94c, Hig94d, Rei89c, Rei89e, Ano93e, Ano97b, Com89, Mux09, Nag09, Rei87e, ANS95]. McFarland [Ano86f]. Meek [BGvE^{+}97]. Meeting [Ano82a, Ano85p, Ano85q, Ano90i, Iwa00, Mei94a, MCL^{+}95, Rei86b, Rei86a, Rei86c, Rei87d, Rei87e, Rei87b, Rei88, Rei89c, Rei89d, Rei89e, Rei90, Rei00b, Ano00e, Mux13, Rei87a, Rei87f, Rei87c, Rei87g]. Meetings [Ano82a, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano90i, Ano90j, Ano91j, Ano91k, Ano92r, Ano92t, Ano92v, Ano92u, Ano92s, Ano93f, Ano94r, Ano94u, Ano95i, Ano95h, Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Clo99, Ano97d, Mux09, Nag09]. Meissner [Ano95b, Sof99]. Members [Mei96c]. Memory [Ste03, Mar03a]. Merged [Ano95a]. Message [Rei99]. messages [Eps96, Eps97]. Metcalf [Sny04, Ano85h, Ano85i, Ano91g, Ano96i, Ano96l, Ano97b, Ano97n]. method [Dec08]. methodology [Kai03a]. Methods [Ano82b, Ano01c, Dup94]. Michael [Sny04, Ano85h, Ano85i, Ano96i, Ano96l, Ano97b, Ano97n]. Microcomputers [Ano85h]. Micros [Ano84e, Ano85e, Mei84a, Mei84b]. Microsoft [Ano95o, Ano97k]. Migrating [Cis94]. Migration [Ano97k]. Milestones [Mei90c, Mei91b]. millennium [Hey98]. Mind [Ano93f]. Missed [Mei96b]. mixed [EHH09]. mode [SWM01]. Model [Ano94d, Ano97d]. modern [Nag05, Nag06]. Module [Ano98b, BGvE^{+}97, MCL^{+}95, OAB^{+}95, Sch01]. Modules [Ano97b, BGvE^{+}97, Bra96, MCL^{+}95, OAB^{+}95, DTF92, Kai03a, Nag06, Ric00, Sch99, SSH^{+}00]. Moss [Mos90]. most [Gre98]. Moving [Ano97b]. MPI [Ano98b]. MPMD [Sie16]. Multiple [Don02]. Muxworthy [Nag09]. myth [Sin99]. myths [Ded99].

N. [Ano96l, Ano01c]. N1122 [ANS95]. NAG [CS07c, CS15c, CS19b, Chi20]. NAGWare [Ano95m, Ano01c]. Names [OAB^{+}95]. Naming [Mei90d]. Nara [Rei00b]. National [Ano87d, Ano92r, Ano92s, Ano94s, Ano94u, Ano95i, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Ano00d, Ano00e, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Com89, Clo99, Ano97d, Ano93f]. NCC [Wag83b]. Needed [Ano92a, Ano92b, Ano93a, Ano94c, Mei91a, Mei92c, OAB^{+}95, Wag82b]. Negative [WCH^{+}03]. Nested [EEV^{+}96]. Net [EEV^{+}96, BGvE^{+}97, BRZ^{+}97, CBFM^{+}98, FRM^{+}97, MCL^{+}95, dPGS^{+}96]. Netherlands [Mux13]. networks [Cur18]. neural [Cur18]. News [Ano87d, Ano06a, Ano06b, Ano06c, Ano17]. Newsletter [Mei98a, Mei98b, Ano85j, Ano84g]. Newsletters [Ano90g, Ano90h, Ano91h, Ano91i, Ano92h, Ano92j, Ano93c, Met99g]. next [Hey98, NR05, Rei06, Rei10]. No [Cla83]. Nominations [Mei98a]. Non [Hen90, Mei96c]. Non-Contiguous [Hen90]. Nonzero [Don02]. Nonzero-Rank [Don02]. North [Ano06b]. Norwegian [Ano86f]. Note [Bus17, Cle09, Mar08b, Mei82a, Mei90b]. Notes [Ano94k, EEV^{+}96]. Notice [Met99g, Met99h]. Notices [Ano84g]. Nov [Rei89c, Ano92r]. November [Ano85p, Mux09, Nag09, Rei87f, Rei87g]. NT [Ano97k]. Number [BRZ^{+}97]. Numerical [Don01a, Don01b, Ano96h, Met94]. NY [Rei89e].

obituary [BGvE^{+}97]. Object

[Ano98a, BGvE^{+}97, Clo94, Coh97, Coh99, Dup94, Her94, Hir96, Lut95, Mar03b, Mei92a, Pat93, Sch98, Ano97d, DNS97, MCC96, Sch93, She00]. object-based [She00]. Object-Oriented [Ano98a, BGvE^{+}97, Clo94, Her94, Mei92a, Lut95, Mar03b, Pat93, Sch98, DNS97, Sch93]. Objects [Jon95, MCL^{+}95, Ano97d]. Obsolete [Mei95b]. Oct [Rei90]. October [Iwa00]. Off [Ano96h]. Official [Ano98b]. Ojbect [Hat03]. Ojbect-Based [Hat03]. Old [Mei87a, Ada86, Ano96k]. On-Line [Ano97h]. OOF [Mei96b]. OOP [Don01a, Sch93]. opaque [Dec08]. OPEN [Ano85j]. OpenCoarrays [FR15]. OpenGL [Ano97a, Ano98b, Mit99]. OpenMP [Kuh98, LP00]. operations [Mar14b, Mei97c]. Optimization [Ano96f, BGvE^{+}97, Hat03]. Optional [Mei93]. Oracle [CS19b, Chi20]. ordered [Mei97c]. ordering [DTSB93]. Organization [Ano93f, Ano94v, Ano94t, Ano94u, Ano95i, Ano95h, Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano98e, Ano98f, Ano98g, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, Clo99, Ano97d]. Organized [Ano94e]. Orientation [Ano83b, Coh97, Coh99]. Oriented [Ano96h, Ano96i, Ano98a, BGvE^{+}97, Clo94, Her94, Mei92a, DNS97, Dup94, Hir96, Lut95, Mar03b, MCC96, Pat93, Sch93, Sch98]. Orleans [Rei88]. Other [Ano82c, Ano83g, Ano84g, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano91e, Ano92p, Ano92q, Ano93d, Ano94n, Ano95k, Ano95n, Ano95m, Ano97h, Ano97i]. Otherwise [BRZ^{+}97]. Oulu [Ano00e, Rei00b]. OUP [Sny04]. Output [Mei93, Mar17, Mei89c, Pag05]. overloaded [Mar14b]. Overview [Koe92]. Oxford [Sny04].

Package [FS01]. Page [EEV^{+}96]. Palo [Rei89d]. Panel [Koe98b]. Papers [Ano83a, Ano99a, Gre98, Met98]. Parallel [Ano94e, Ano97h, CS05b, Kuh98, Lon16, BCS92, Cur18, NR98, Rei92, Ano94w, Ano95n]. parallelisation [TM12]. parallelism [Bus10]. Parallelization [LP00]. Paris [Ano82b, Rei89c]. Park [Rei90]. Parlez [Ano96h]. Parlez-vous [Ano96h]. Part [DN99, Don02, Hig94a, Met02b, Met02c, Met03b, DNS99, DN00a, Ano93e, Coh97, Coh99, Don01a, Don01b, Hig94b, Hig94c, Hig94d, Mei90e, Met99c, Met99d, Met99e, Met00c, Met00a, Met00b, Met01c, Met01d, Met01b]. Partial [Mei89c]. Partial-record [Mei89c]. Participants [MCL^{+}95]. parts [Smi85]. Pass [MCL^{+}95]. passing [Dec08]. Past [Met99h, Cla82, Cla83]. patterns [GD06, Mar06, Mar08a]. pay [Mei95a]. pays [Mei95a]. PC [Ano85m, Ano95m, Ano96i, Ano97k]. PCF [The91, Rei92, WRV^{+}92]. PDTR [ISO96b]. Performance [Ano92k, Ano92v, Ano93e, Ano93f, Ano94v, Ano98d, BGvE^{+}97, Iwa00, Koe92, MCL^{+}95, Pri99]. Period [Ano87d]. Periodicals [Ano90f, Ano90e, Ano91f, Ano91g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano93c, Ano94i, Ano94j, Ano94k, Ano95k, Ano96h, Ano96i]. Permit [Ano95a]. Personal [Wag83b, Gui87, Mux09]. PFortran [BCS92]. PGI [Ano06c]. Philips [Ano95b]. Philosophy [Ano96k]. photo [Wag03]. Phrase [Ros95]. physics [Pre95]. PL22.3 [Mux13]. Placement [EEV^{+}96]. Plans [Ano94v, Ano94u, Ano94x]. platform [Sin99]. Plus [Ano96l]. plusFort [Ano95o]. Point [OAB^{+}95, ISO96a]. Pointer [BGvE^{+}97, Bur87, SM87, Mei88a, Rol95]. Pointers [BGvE^{+}97, Mar17, OAB^{+}95, Ano97d, Ble96, Cle09, Dec08, Mar92, Mei97c]. poll [Bur09]. Polymorphic [FS01, Ano97d, LC11b]. Polymorphism [Bad07a, Bad06]. Polytron [Ano86f]. Poor [Sny97]. Port [Rei89e]. Portability [Ano83a, Ano85j, Nag99, Sin99]. portable [Vow99]. POSIX [Ano89k]. possible [TM12]. Postscript [Met99f]. Power [Ano95o, Ano96i]. Power8E [FR15]. PowerStation [Ano97k]. practitioner [Pat93]. Precision [Ano98b, Sch01]. preliminary [Mei96a]. preprocessor [Ano96k]. Preprocessors [Ano95n, OAB^{+}95]. prescriptive [RW99b]. Present [Ano83b, Cla82, Cla83]. presented [Ano00e]. Press [AFI84, Ano96h, Mar03b]. Pretty [Met02e]. Prettyprinters [Ano95m]. print [BGvE^{+}97]. Printing [Met02e]. problems [Ber85b, Pri99]. Procedure [OAB^{+}95, Cle09, Pag96]. Procedures [Hig94b, Hig94c, Mei95b, OAB^{+}95, Wag95, Gor08, Gor14a, Gor14b, Vow99]. Proceedings [Ano94a]. Processing [Ano87d]. Processor [Ano97k]. Product [Ano82d, Ano83h, Ano95o, Ano99d, Ano99e]. Products [Ano97n, Ano97l, Ano97m, Ano98i]. Program [Ngu03, LP00, Met06, Ano01b, Sie16]. Programmer [Ano91g, Ano95o]. Programming [Ano87d, Ano94d, Ano96h, Com89, EEV^{+}96, Hat03, Hol95, Mar03b, Mei82b, Met20, CS05b, EHH09, Hir96, Mar01, Mar19, McC05, MCC96, Met98, NR98, Sch98, Wag84, Ano95b]. Programs [Ano85i, Sny00, CS06a, Gui87]. Progress [Ano97b, Wag93]. project [Nag99, Ano84h]. Promotional [Mei83b]. Proposal [Ano98a, Her94]. proposals [Mei88a]. Proposed [Ano87d, Rag82, Wag82b]. Proposes [Mei92c]. PROTRAN [Ano86f]. Provides [Ano98b]. Public [Ano87d, Ano90b, Mei89b, Mei89a, Mei90c, Ano01c]. Publications [Ano84g, Ano85g, Ano85h, Ano85i, Ano85j, Ano86c, Ano86d, Ano87e, Ano87f, Ano87g, Ano88b, Ano89d, Ano89e, Ano89f, Ano90f, Ano90e, Ano91f, Ano91g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano93c, Ano94h, Ano94i, Ano94j, Ano94k, Ano95j, Ano95k, Ano96g, Ano96h, Ano97g, Ano97h, Ano97i, Ano98h]. Published [Ano97b, Met99h]. Pure [FRM^{+}97, Wag95]. PyF95 [CL10, LC11a].

Quality [MDB93]. Quantum [Mar04b]. Question [Ano84j]. Questionnaire [Ano83b, Ano01e, Met00d, Met01e]. Questions [Ano83b, Mei83b, Mar03c, Tay97]. Quotes [Ano85k, Wag82b].

Random [BRZ^{+}97]. Rank [Don02]. Rates [Mei92c]. Rationale [Don07]. RATIONALES [Ded95]. Re [Bro92, DTSB93, OAB^{+}95]. Re-ordering [DTSB93]. read [BGvE^{+}97]. Readability [Ano83a]. Readable [Ano91g]. Readers [Ano92a, Ano92b, Ano93a, Ano94c, Mei91a, Mei92c]. Real [Ano94y, Ano94v]. Real-Time [Ano94y, Ano94v]. REALLOCATE [Ano95l]. really [Koe98a]. Recipes [Ano96h, Met94]. Reconsidered [Red93]. record [Mei89c]. Recoup [SVM^{+}93]. Red [Mei91a]. refactoring [OFKJ10]. reference [Car10, LC11b]. References [Don02]. Reflections [Wag97a, Wag97b, Wag97c]. Regularization [ISO96b]. Reid [Sny04, Ano85n, Ano91g, Rei86b, Rei86a, Rei86c, Rei87a, Rei87f, Rei87c, Rei87d, Rei87e, Rei87b, Rei88, Rei89a, Rei89c, Rei89d, Rei89e, Rei90]. Rejection [Mei92c]. Related [Ano84g, Ano92p, Ano92q, Ano93d, Ano94n]. Release [AFI84, Ano87d, Ano01c]. reliability [Ber85b, Mei85d]. remarks [Ano99c]. Replacement [BGvE^{+}97]. reply [Nag09]. Report [Ano85i, Ano85j, Ano85p, Ano85q, Ano90i, ISO96b, Iwa00, NR98, WRV^{+}92, ISO96a, Mux13, Ano00e]. Reports [Ano82c, Ano85n, Ano92h, Ano92l, Ano93c, Ano94k, Ano95a, Ano98b, Mei96d, Mos90, Rei86b, Rei86a, Rei86c, Rei87d, Rei87e, Rei87b, Rei88, Rei89c, Rei89d, Rei89e, Rei90, Ano85j, Rei87a, Rei87f, Rei87c]. Request [Ano82e]. Requirements [Ano84h, Mei96b]. research [Ano85j, LP00, Ano94d]. Resolution [Ano97d]. Resolutions [Ano85n]. Resources [Ano94o, Ano95g, Ano95n, Ano95m, Ano96l, Ano97n, Ano97l, Ano97m, Ano98i, Mei94c, CS06b, Met95]. response [Rei03]. Restricted [OAB^{+}95]. Restructures [Ano95o]. Results [BGvE^{+}97, ISO96b]. retire [Ano06b]. Returning [Vow02]. rev [Ano83b]. Review [Ano87d, Cis94, Mar03b, Mar14a, Mei89b, Mei89a, Mei90c, Sny04, Met20]. Reviews [Ano82c, Ano90f, Ano91g, Ano95b, Ano01d, Mar09]. Revised [Ano87d, Mei90e]. Revision [Com89, CS10a, CS12b, CS15b, CS15a, CS16a, CS16b, CS16c, CS17a, CS17b, CS18b, CS18c, CS18a, CS19a, Chi20, AR87, CS11a, CS11b, CS11c, CS12c, CS12a, CS13b, CS13c, CS13a, CS14, Rei06, Wag84]. Revisited [SG02, Ste03]. Revolution [Mei90c, Mei91b]. Richardson [Rei90]. RM/Fortran [Ano86f]. Role [Ell98, Mei89b]. Rouson [Mar14a]. Routines [Don01a, Don01b]. Rules [Rag82]. Rumor [Mei84b]. Ryan [Ano86f]. Ryan-McFarland [Ano86f].

S8 [Com89, Miy87]. Saga [Mee90]. Salford [Ano95o]. Same [CBFM^{+}98]. sans [SM87]. SAVE [CBFM^{+}98, BGvE^{+}97]. SC22 [Ano00e, Rei00b]. SC33 [ANS95]. SC33/WG5 [ANS95]. SC33/WG5-N1122 [ANS95]. Schedule [Mei94a, Mei94c]. Scheduled [Ano97b]. Science [Ano97k, Ano01c, Ano00a]. Science-Engineering [Ano97k]. Scientific [Bad07b, MDB93, Ano96l, Mar14a]. scripting [Mar02]. Seattle [Rei87e]. Second [Mei89b, Mei89a, Mar05]. semantics [RW99b]. Semiautomatic [Kai03b]. Sep [Rei89c]. separate [Ano83a]. Separating [Ano97b]. September [Rei00b, Ano85n]. Sequence [Hig94d]. Sequencing [Ngu03]. service [Bus10]. Services [Ano96l]. Session [EEV^{+}96, Koe98b]. Sets [Mei96b]. SHARE [Ano84h]. Shares [Ano94d]. Short [BGvE^{+}97, Mei92a, Wag83a]. Should [Jon95, Mei93, Mei95b, Mei95a]. Side [FRM^{+}97]. SIG [Met99g]. SIGForth [Ano95a]. Significant [BGvE^{+}97, Ber85b]. SIGNUM [Ano84g]. SIGPLAN [Ano84g, Ano94a, Ano95a]. simply [Arm12]. Simulates [Ano97h]. Sites [Ano98b]. Snobs [BGvE^{+}97]. Society [Ano85j]. Softool [Ano85m]. Software [Ano83a, Ano83i, Ano85f, Ano85h, Ano85l, Ano85m, Ano86e, Ano86f, Ano87h, Ano87i, Ano87j, Ano88c, Ano89g, Ano89h, Ano89i, Ano90g, Ano90h, Ano91d, Ano91e, Ano91h, Ano91i, Ano92h, Ano92j, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano93c, Ano93d, Ano94o, Ano94n, Ano94m, Ano94l, Ano95n, Ano95m, Ano95o, Ano95p, Ano95q, Ano96i, Ano96l, Ano96j, Ano96k, Ano97n, Ano97l, Ano97m, Ano97j, Ano97k, Ano98i, Ano01b, Ano01c, Ano06c, Hou05, Mar14a, Mei92a, MDB93, NAG85, Smi85, SNB^{+}98, Lut95, Sin99, TM12, Ano85j, Ano96l]. Solver [Don01b]. Solving [Mei97b]. Some [OAB^{+}95, Rei89b, Tay97, Rei89a, Sch93]. SOS [Ano83a]. Source [Mei93, Met91, Mei90b, TM12]. Speak [Ano93f, Mei87b]. Speakers [Ell98]. Special [Ano82c, Ano93e, Hig94b, Hig94c, Hig94d, Mei83b, Mei92e, Mei96d, NR98, The91, Wag93, ANS95]. Specification [Ano93e, Ano93f, Ano94v, Ano97b, Hig94a, Hig94b, Hig94c, Hig94d, Mei96a]. specifications [OFKJ10]. Spotlight [Ano95o, Ano95p, Ano97j, Ano97k, SNB^{+}98, Ano96j, Ano96k]. Spotlights [Ano95q]. Standard [Ano87d, Com89, OAB^{+}95, Run95, AR87, CS07a, CS07b, CS08a, CS08b, CS09c, Lah87, Met87, Rei10, Wag83a, Ano83a, Ano83b, Ano92r, Ano93f, Ano94p, Ano94v, Ano94u, Ano94y, Ano95n, Ano02c, Ano03b, Cla82, Cla83, Clo99, EEV^{+}96, FRM^{+}97, Mei83b, Mei92c, Mei92b, Mei92d, NR05, SWM01]. Standardization [Ano84j, OAB^{+}95]. Standardize [BGvE^{+}97]. Standardized [EHH09]. Standards [Ano82a, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano90i, Ano90j, Ano91j, Ano91k, Ano92r, Ano92t, Ano92v, Ano92u, Ano92s, Ano93f, Ano94p, Ano94r, Ano94q, Ano94s, Ano94u, Ano94x, Ano94w, Ano94y, Ano02d, Ano02c, Ano03b, CS10a, CS12b, CS15b, CS15a, CS16a, CS16b, CS16c, CS17a, CS17b, CS18b, CS18c, CS18a, CS19a, Met03b, Rei00b, Rei02b, CS09b, CS09a, CS10b, CS10c, CS11a, CS11b, CS11c, CS12c, CS12a, CS13b, CS13a, CS14, Gor08, Gor14a, Gor14b, Mei90d, Rei01, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano89l, Ano91j, Ano91k, Ano93f, Ano94v, Ano94t]. Standards [Ano94u, Ano95a, Ano95i, Ano95h, Ano95r, Ano96c, Ano96d, Ano96e, Ano97b, Ano97e, Ano97f, Ano98a, Ano98e, Ano98f, Ano98g, Ano99b, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano02d, Ano02c, Ano03b, CS13c, CS20, Clo99, Ell98, Mei95b, Mei95a, Ano97d]. Statements [BGvE^{+}97, Gor08, Gor14a, Gor14b]. Station [Ano95o]. Status [Ano85j, BGvE^{+}97, Wag84]. Storage [Hig94d, MCL^{+}95]. StratCom [Ano86f]. Stream [Pag05]. strengths [Koe98b]. String [MCL^{+}95, Vow02, Gil96]. strings [Sch92, Bro92]. Structure [Don02, Ros95, Mai96, Pag96]. Structured [Ano85j]. structures [OAB^{+}95]. students [Pre95]. Studio [Ano97k, CS09d]. Subcommittee [Wag82b]. Submissions [ACM94]. submodules [Bad06]. Subscription [Mei92c]. Subscriptions [Mei96c]. Subset [Ano95p, Ano96l, FRM^{+}97, Run95, BGvE^{+}97]. Subsets [OAB^{+}95, Mei97a]. Succeeding [OAB^{+}95]. Successor [Ano95q]. Sudoku [Met06]. suggestions [Tay97]. Suite [Ano95a, Ano96l, NAG85]. Summary [AW89, Ano96j, Mei89d, Mei90e, NAG85]. Summation [EEV^{+}96]. Sun [Ano85m, CS09d]. sup [Ano83b]. supercomputer [Ano85j]. SUPERNET [Ano85j]. Support [Ano98b, CS15b, CS15a, CS16a, CS16b, CS16c, CS17a, CS17b, CS18b, CS18c, CS18a, CS19a, CS20, FR15, CS07a, CS07b, CS08a, CS08b, CS09b, CS09c, CS09a, CS10a, CS10b, CS10c, CS11a, CS11b, CS11c, CS12b, CS12c, CS12a, CS13b, CS13c, CS13a, CS14]. Survey [Val83, Wag83b]. Switch [EEV^{+}96]. switches [CS06a]. Symbol [EEV^{+}96]. syntax [Gil96]. Systems [Ano87d, Com89, DTSB93, Mei97b, Ano85i, Ano86f].

Tagged [Ano97d]. take [Eps97]. Target [Wag95, Rol95]. Targets [BGvE^{+}97]. TC8 [Ano83b]. Teach [BGvE^{+}97]. Technical [AWP82, Ano85j, Ano92h, Ano95a, Ano98b, ISO96b, ISO96a, Mei96d, Gre98]. Technically [Mei91a]. technology [Ano85j, Ano86f]. templates [Sny97]. templating [CL10, LC11a]. Ten [Ano00c]. Test [Ano95a, Ano01c]. Testing [Ano92p, Ano92q, Ano93d, LC11a]. Tests [Ano95n]. Text [Ano91g, Ano92s]. Textbook [Ano96h, Ano98b, MCL^{+}95, Ano85j]. Textbooks [Ano96h, Ano82c]. Thank [Mei90d]. them [Ber85b]. theme [Cle99]. There [Ell98, Del00b]. Thin [Mei99]. Third [Mei90c]. Thoughts [Bur87, Mei92d, Sch93]. threads [OAB^{+}95]. Time [Ano87d, Ano94y, Vau00, Pol13, Ano94v]. Times [Mei82a]. Tokyo [Iwa00]. too [Wag85b, Wag85a]. Tool [Ano83a, Ano96l, Ano01b, ALN18]. Toolkit [Ano95o]. TOOLPACK [NAG85, Ano95m]. TOOLPACK/1 [NAG85]. Tools [Ano82b, Ano86f, Ano94e, Ano94w, Ano95n, Ano95m, Ano97j, Goe08, Met02b, Met02e, NAG85, Hou05, Met00b, Nag05]. toolset [LC11a]. Top [Ano00c]. touches [Mei90d]. traditional [Lut95]. Train [FRM^{+}97, Mei96b]. Trajectory [Hat03]. Triangular [MCL^{+}95]. Trivia [Lau92, Mar03c]. Truly [Vow01]. TS29113 [CS18a, CS19a, CS18b]. Turn [Mei86]. Turn-Around [Mei86]. TV [Ano90b]. Two [CBFM^{+}98, Eps97]. TX [Rei90]. Type [Ano97d, ISO96b, SM87]. Types [Wal00, Gil01, Gor14a, Gor14b, Mar03a, Ste03].

UCBTest [Ano95m]. UK [Ano00e]. Undefined [BGvE^{+}97]. Unicomp [Ano95a]. unit [LC11a]. unit-testing [LC11a]. University [Mar03b]. UNIX [Mei84b]. Update [Ano92r, Mei92e, TM12]. Upper [FRM^{+}97]. Urbana [Rei89c]. Usage [Ano83b]. Use [Del00a, Wal00, Del00b, Ric00]. User [Ano97j, Don01a, Don01b, EEV^{+}96, Iwa00, Mar14b]. User-defined [Mar14b]. Users [Ano83b]. Uses [CBFM^{+}98, Sny01]. Using [Cla82, Cla83, Rei93a, DTF92, Dup94, LP00, Sch99, Sny00]. utilities [Ano95n, Met99d]. utilization [RW99a]. utilizing [LC11a].

v2.0 [Ano95o]. Validation [Ano92p, Ano92q, Ano93d]. valued [Mar08b]. Values [MCL^{+}95]. Variable [Ano98b, Sch01, BGvE^{+}97]. Variables [BGvE^{+}97, CBFM^{+}98, Mei95b, MCL^{+}95, OAB^{+}95, Ano97d]. variation [Cle99]. Varying [MCL^{+}95, Sch92, Vow02, Bro92]. VAX [Kha91]. VAX-11 [Kha91]. Verifications [Ngu03]. versa [She99]. Version [Ano93e, CS20, Miy87, Ano01c, Hig94a, Hig94b, Hig94c, Hig94d]. Versions [Ano98b]. via [EEV^{+}96, Mar03b]. vice [She99]. Vienna [Rei90]. View [Mei82b, Bus10, Mux09, Pat93]. Visibility [BGvE^{+}97, OAB^{+}95]. Visual [Ano01b, Ano97k]. VisualFortran [Ano97j]. Vol. [Cla83]. Volatile [BGvE^{+}97]. Volunteers [Wag82b]. Vote [Mei84b, Mei96b, WCH^{+}03]. vous [Ano96h]. vs [Ada86, Ano96f, Mar92, Mei87a, Rei93b].

Wants [MCL^{+}95]. Watfor [Red91]. Watfor-77 [Red91]. weaknesses [Koe98b]. Welcomes [MCL^{+}95]. Well [MCL^{+}95]. West [Mei98b]. WG5 [Ano00e, Rei89c, Rei00b, Rei01, Rei02b, Ano90i, Ano99c, Mei96b, Mux09, Mux13, Nag09, Rei90, Rei99, Rei00b, Rei03, Wag95]. WG5-N1122 [ANS95]. Where [Mei88b, MCL^{+}95]. WHILE [Red93, Rei93b, Rei93a]. Whither [Mei83c, Mei84b, MR91, Nag02]. Who [CBFM^{+}98, Mei95a, Mei95a]. will [Koe98a, Mei91b]. Windows [Ano95o, Ano97k]. without [MCL^{+}95]. Word [Ano85o]. work [Wag84]. Working [Ano92r, Ano92s, Ano93f, Ano94s, Ano94u, Ano95i, Ano98b, Mei89b, Mei98b, Miy87, Wag95]. Workshop [Ano92r, Kuh98]. Workstations [Ano85m]. world [Gre98]. Writing [She99, She00, Vow99]. WWW [EEV^{+}96]. WY [Rei89c].

X [Ano85j]. X/OPEN [Ano85j]. X3 [Ano87d]. X3.9 [Ano87d, Com89, Com89]. X3.9-1978 [Com89]. X3.9-198X [Com89, Ano87d]. X3J14 [Ano89l]. X3J3 [Ano82a, Miy87, Rei89c, AWP82, Ano82a, Ano83j, Ano83k, Ano83l, Ano83m, Ano84i, Ano84j, Ano85p, Ano85q, Ano85n, Ano85o, Ano86g, Ano86h, Ano87k, Ano87l, Ano87m, Ano88d, Ano89l, Ano89k, Ano89j, Ano90i, Mei89e, Rei86b, Rei86a, Rei86c, Rei87a, Rei87f, Rei87c, Rei87d, Rei87e, Rei87b, Rei87g, Rei88, Rei89c, Rei89d, Rei89e, Rei90, Wag95]. X3J3/95 [ANS95]. X3J3/95-007r1 [ANS95]. X3J3/S8 [Miy87]. Xia [Mar14a]. Xiaofeng [Mar14a]. XML [Mar04a, Mar05]. Xu [Mar14a].

Year [Koe98a]. York [Sny04]. yourself [Ano96h].

ZeroOne [Ano85j]. Zeus [Ano06c].

References

ACM:1994:AAG

[ACM94] ACM. ACM announcement:

General guidelines for ACM electronic submissions. ACM Fortran Forum, 13(3):21, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1982:C

[Ada82] Jeanne Adams. Correspondence.

ACM ForTec Forum, 3(2):44, June 1982. CODEN ???? ISSN 0735-3731.

Adams:1986:GFV

[Ada86] Jeanne C. Adams. The glorious future vs. the good old days (editorial). ACM Fortran Forum, 5(1):1, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

AFIPS:1984:PRF

[AFI84] AFIPS. Press release: Fortran celebrated in Annals of the History of Computing. ACM ForTec Forum, 3(1):9–10, March 1984. CODEN ???? ISSN 0735-3731.

Aharonian:1985:AFC

[Aha85a] Gregory Aharonian. Automated Fortran conversion. ACM Fortran Forum, 4(4):13–19, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Aharonian:1985:LKF

[Aha85b] Gregory Aharonian. Let’s kill future Fortrans. ACM Fortran Forum, 4(3):21–23, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anderson:2018:HAT

[ALN18] Michelle Anderson, Sisi Garnet Liu, and Daniel Nagle. The h2m-AutoFortran tool for facilitating C–Fortran interoperability. ACM Fortran Forum, 37 (2):8–18, August 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1982:AFS

[Ano82a] Anonymous. ANSI Fortran Standards Committee (X3J3) — actions at recent meetings: X3J3 82nd Meeting, May 1982. ACM ForTec Forum, 1(1):15, July 1982. CODEN ???? ISSN 0735-3731.

Anonymous:1982:ICT

[Ano82b] Anonymous. International Conference on Tools, Methods, and Languages, May 17–19, 1983, Paris, France. ACM ForTec Forum, 1(2):12, December 1982. CODEN ???? ISSN 0735-3731.

Anonymous:1982:LFT

[Ano82c] Anonymous. Literature: Fortran 77 textbooks; reviews of Fortran 77 textbooks; other reviews; reports; special issue of the Annals to feature history of Fortran. ACM ForTec Forum, 1(2):9–10, December 1982. CODEN ???? ISSN 0735-3731.

Anonymous:1982:PA

[Ano82d] Anonymous. Product announcements. ACM ForTec Forum, 1 (2):11, December 1982. CODEN ???? ISSN 0735-3731.

Anonymous:1982:RI

[Ano82e] Anonymous. Request for information. ACM ForTec Forum, 1 (2):12, December 1982. CODEN ???? ISSN 0735-3731.

Anonymous:1983:ALA

[Ano83a] Anonymous. Announcements:

Learn about Fortran 77, Fortran 8X, or both; international failure journal SOS; Fortran 77 portability and readability guidelines; computer conference in China: Call for papers; software tool for management and documentation; what’s this about a separate British Fortran Standard? ACM ForTec Forum, 2(2):8–9, June 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:ETS

[Ano83b] Anonymous. ECMA/TC8/83/3 (sup 82/18) rev LPM Aug 83: Questionnaire: Present usage of the Fortran language and desired orientation of the future Standard Fortran 8X: Questions to the users. ACM ForTec Forum, 2 (3):11–15, September 1983. CO-DEN ???? ISSN 0735-3731.

Anonymous:1983:FCa

[Ano83c] Anonymous. Forthcoming conferences. ACM ForTec Forum, 2 (1):7, March 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:FCb

[Ano83d] Anonymous. Forthcoming conferences. ACM ForTec Forum, 2 (4):9, December 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:La

[Ano83e] Anonymous. Literature. ACM ForTec Forum, 2(1):7, March 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:Lb

[Ano83f] Anonymous. Literature. ACM ForTec Forum, 2(4):8, December 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:OA

[Ano83g] Anonymous. Other announcements. ACM ForTec Forum, 2 (1):8, March 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:PA

[Ano83h] Anonymous. Product announcements. ACM ForTec Forum, 2 (4):9, December 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:S

[Ano83i] Anonymous. Software. ACM ForTec Forum, 2(1):8, March 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:SAAa

[Ano83j] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3. ACM ForTec Forum, 2(1):2–3, March 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:SAAb

[Ano83k] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3. ACM ForTec Forum, 2(2):2–3, June 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:SAAc

[Ano83l] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3. ACM ForTec Forum, 2(3):2–4, September 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1983:SAAd

[Ano83m] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3. ACM ForTec Forum, 2(4):2–3, December 1983. CODEN ???? ISSN 0735-3731.

Anonymous:1984:A

[Ano84a] Anonymous. Announcements.

ACM Fortran Forum, 3(3):13, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1984:CFC

[Ano84b] Anonymous. Certified Fortran compilers. ACM ForTec Forum, 3(1):12, March 1984. CODEN ???? ISSN 0735-3731.

Anonymous:1984:FCa

[Ano84c] Anonymous. Forthcoming conferences. ACM ForTec Forum, 3 (1):8, March 1984. CODEN ???? ISSN 0735-3731.

Anonymous:1984:FCb

[Ano84d] Anonymous. Forthcoming conferences. ACM Fortran Forum, 3 (3):11, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1984:FM

[Ano84e] Anonymous. Fortran 77 on micros. ACM Fortran Forum, 3(3): 14–15, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1984:L

[Ano84f] Anonymous. Literature. ACM ForTec Forum, 3(1):7, March 1984. CODEN ???? ISSN 0735-3731.

Anonymous:1984:PNC

[Ano84g] Anonymous. Publications: New Computer Language magazine; Fortran related articles in recent issues of SIGPLAN Notices; Fortran related articles in recent issues of SIGNUM Newsletter; other recent Fortran related articles. ACM Fortran Forum, 3(3): 11–12, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1984:SFP

[Ano84h] Anonymous. SHARE Fortran Project requirements. ACM ForTec Forum, 3(1):11, March 1984. CODEN ???? ISSN 0735-3731.

Anonymous:1984:SAAa

[Ano84i] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; other Standards activities. ACM ForTec Forum, 3(1):3–6, March 1984. CODEN ???? ISSN 0735-3731.

Anonymous:1984:SAAb

[Ano84j] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; other Standards activities; recent changes to Fortran 8X; a Fortran 77 question; more comments on Fortran standardization. ACM Fortran Forum, 3(3): 2–10, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:FCa

[Ano85a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 4(1):16, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:FCb

[Ano85b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 4(2):10, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:FCc

[Ano85c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 4(3):13, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:FCd

[Ano85d] Anonymous. Forthcoming conferences. ACM Fortran Forum, 4 (4):7, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:FMF

[Ano85e] Anonymous. Fortran 77 on micros (follow-up). ACM Fortran Forum, 4(1):16, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:HS

[Ano85f] Anonymous. Hardware and software. ACM Fortran Forum, 4 (2):11, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:P

[Ano85g] Anonymous. Publications. ACM Fortran Forum, 4(4):7, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:PEF

[Ano85h] Anonymous. Publications: Effective Fortran 77, by Michael Metcalf; software catalog: Microcomputers. ACM Fortran Forum, 4(2):10, June 1985. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:PNJ

[Ano85i] Anonymous. Publications: New journal: Future Generations Computer Systems; book: Effective Fortran 77, by Michael Metcalf; report: Design conventions for Fortran programs, by Michael Metcalf. ACM Fortran Forum, 4(1):16, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:PSM

[Ano85j] Anonymous. Publications: Software Magazine (IEEE Computer Society); ZeroOne SUPER-NET (supercomputer newsletter); technical reports from Argonne; Structured Fortran for Business (textbook); X/OPEN portability guide (common applications environment); Esprit ’84 status report (European technology research). ACM Fortran Forum, 4(3):14, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:Q

[Ano85k] Anonymous. Quotes. ACM Fortran Forum, 4(3):16, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:S

[Ano85l] Anonymous. Software. ACM Fortran Forum, 4(3):15, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:SFI

[Ano85m] Anonymous. Software: Fortran 77 for IBM PC; Softool Fortran environment on Sun workstations. ACM Fortran Forum, 4 (4):8, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:SAAc

[Ano85n] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3: International meetings; other Standards activities; John Reid reports (X3J3 meetings, July and September 1985); more about the composite Do construct; international Fortran resolutions. ACM Fortran Forum, 4(3):2–12, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:SAAd

[Ano85o] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings; other Standards activities; latest word on the Fortran 8x composite Do construct. ACM Fortran Forum, 4(4):2–6, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:SAAa

[Ano85p] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; other Standards activities; report of X3J3 meeting 92, 12–16 November 1984; report of X3J3 meeting 93, 11–15 February 1985; letter from Jeanne Adams to John Dvorak. ACM Fortran Forum, 4 (1):2–15, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1985:SAAb

[Ano85q] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3: Report of X3J3 meeting, 92, 6–10 May 1985 and other Standards activities. ACM Fortran Forum, 4 (2):2–9, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:FCa

[Ano86a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 5(1):13, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:FCb

[Ano86b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 5(2):9, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:Pa

[Ano86c] Anonymous. Publications. ACM Fortran Forum, 5(1):11, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:Pb

[Ano86d] Anonymous. Publications. ACM Fortran Forum, 5(2):9, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:S

[Ano86e] Anonymous. Software. ACM Fortran Forum, 5(2):9, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:SFN

[Ano86f] Anonymous. Software: FORTVER (Norwegian Institute of Technology; Infotron Limited); PRO-TRAN (IMSL); RM/Fortran (Ryan-McFarland Corporation); Fortran tools (StratCom Systems; Polytron Corporation). ACM Fortran Forum, 5(1):12, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:SAAa

[Ano86g] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings; other Standards activities. ACM Fortran Forum, 5(1):2–3, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:SAAb

[Ano86h] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3: International meetings; other Standards activities. ACM Fortran Forum, 5(2):2–3, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:FCa

[Ano87a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 6(1):15, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:FCb

[Ano87b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 6(2):21, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:FCc

[Ano87c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 6 (3):7, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:NRX

[Ano87d] Anonymous. News release (X3, information processing systems): The time has come: X3 announces public review and comment period on Draft Proposed Revised American National Standard X3.9-198x, Programming Language Fortran. ACM Fortran Forum, 6(2):1, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Pa

[Ano87e] Anonymous. Publications. ACM Fortran Forum, 6(1):15, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Pb

[Ano87f] Anonymous. Publications. ACM Fortran Forum, 6(2):21, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Pc

[Ano87g] Anonymous. Publications. ACM Fortran Forum, 6(3):8, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Sa

[Ano87h] Anonymous. Software. ACM Fortran Forum, 6(1):16–17, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Sb

[Ano87i] Anonymous. Software. ACM Fortran Forum, 6(2):22, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:Sc

[Ano87j] Anonymous. Software. ACM Fortran Forum, 6(3):7, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:SAAa

[Ano87k] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings. ACM Fortran Forum, 6(1):2–3, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:SAAb

[Ano87l] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings. ACM Fortran Forum, 6(2):4–5, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1987:SAAc

[Ano87m] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings. ACM Fortran Forum, 6(3):2–3, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1988:FC

[Ano88a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 7(1):6, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1988:P

[Ano88b] Anonymous. Publications. ACM Fortran Forum, 7(1):7, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1988:S

[Ano88c] Anonymous. Software. ACM Fortran Forum, 7(1):8, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1988:SAA

[Ano88d] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings. ACM Fortran Forum, 7(1):2–3, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:FCa

[Ano89a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 8(1):18, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:FCb

[Ano89b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 8 (2):9, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:FCc

[Ano89c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 8(3):8, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Pa

[Ano89d] Anonymous. Publications. ACM Fortran Forum, 8(1):20, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Pb

[Ano89e] Anonymous. Publications. ACM Fortran Forum, 8(2):11, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Pc

[Ano89f] Anonymous. Publications. ACM Fortran Forum, 8(3):9, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Sa

[Ano89g] Anonymous. Software. ACM Fortran Forum, 8(1):19, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Sb

[Ano89h] Anonymous. Software. ACM Fortran Forum, 8(2):10, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:Sc

[Ano89i] Anonymous. Software. ACM Fortran Forum, 8(3):10, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:SAAc

[Ano89j] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings. ACM Fortran Forum, 8(3):3–4, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:SAAb

[Ano89k] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings; new IEEE Standards Group for POSIX binding to Fortran. ACM Fortran Forum, 8(2):2, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1989:SAAa

[Ano89l] Anonymous. Standards: Activities of the ANSI Fortran Standards Committee X3J3; international meetings; new Standards group X3J14 for Forth language. ACM Fortran Forum, 8(1):4–5, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:AAAb

[Ano90a] Anonymous. ACM announcements: ACM awards. ACM Fortran Forum, 9(2):6, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:AAAa

[Ano90b] Anonymous. ACM announcements: ACM awards; computer history on public TV. ACM Fortran Forum, 9(1):15, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:FCa

[Ano90c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 9(1):16, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:FCb

[Ano90d] Anonymous. Forthcoming conferences. ACM Fortran Forum, 9(2):7–8, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:PAPb

[Ano90e] Anonymous. Publications: Articles, periodicals, books. ACM Fortran Forum, 9(2):9–10, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:PAPa

[Ano90f] Anonymous. Publications: Articles, periodicals, books, book reviews. ACM Fortran Forum, 9 (1):17–19, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:SSNa

[Ano90g] Anonymous. Software: Software newsletters. ACM Fortran Forum, 9(1):20, March 1990. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:SSNb

[Ano90h] Anonymous. Software: Software newsletters. ACM Fortran Forum, 9(2):11–12, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:SIF

[Ano90i] Anonymous. Standards: Information; future meetings; international meetings; report of X3J3 meeting 115, April–May 1990; report of WG5 and X3J3 meetings, August 1990; the future. ACM Fortran Forum, 9(2):2–5, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1990:SII

[Ano90j] Anonymous. Standards: Information: International meetings. ACM Fortran Forum, 9 (1):4–5, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:AAA

[Ano91a] Anonymous. ACM announcements: ACM awards. ACM Fortran Forum, 10(1):5, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:FCa

[Ano91b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 10(1):6–7, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:FCb

[Ano91c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 10(2):5–6, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:FS

[Ano91d] Anonymous. Fortran software. ACM Fortran Forum, 10(2):7– 10, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:OS

[Ano91e] Anonymous. Other software. ACM Fortran Forum, 10(2): 10, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:PAP

[Ano91f] Anonymous. Publications: Articles, periodicals, books. ACM Fortran Forum, 10(1):10–14, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:PMR

[Ano91g] Anonymous. Publications: Machine readable final Fortran 90 text, articles, periodicals; books, book reviews: Brainerd, Goldberg, and Adams: Programmer’s Guide to Fortran 90, Metcalf and Reid: Fortran 90 Explained. ACM Fortran Forum, 10(2):11– 13, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:SN

[Ano91h] Anonymous. Software newsletters. ACM Fortran Forum, 10 (2):10, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:SSN

[Ano91i] Anonymous. Software: Software newsletters. ACM Fortran Forum, 10(1):8, March 1991. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:SIFa

[Ano91j] Anonymous. Standards: Information; future meetings; international meetings; current Standards activities. ACM Fortran Forum, 10(1):2–4, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1991:SIFb

[Ano91k] Anonymous. Standards: Information; future meetings; international meetings; current Standards activities. ACM Fortran Forum, 10(2):2–4, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:AIRa

[Ano92a] Anonymous. Announcements: Input from readers needed. ACM Fortran Forum, 11(2):14, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:AIRb

[Ano92b] Anonymous. Announcements: Input from readers needed. ACM Fortran Forum, 11(3):10, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:FCa

[Ano92c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 11(1):5, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:FCb

[Ano92d] Anonymous. Forthcoming conferences. ACM Fortran Forum, 11(2):12–13, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:FCc

[Ano92e] Anonymous. Forthcoming conferences. ACM Fortran Forum, 11(3):9–10, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:FCd

[Ano92f] Anonymous. Forthcoming conferences. ACM Fortran Forum, 11(4):8, December 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:FCe

[Ano92g] Anonymous. Forthcoming conferences. ACM Fortran Forum, 12(2):11, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:PBA

[Ano92h] Anonymous. Publications:

Books, articles and technical reports, periodicals, software newsletters. ACM Fortran Forum, 12(2):9–10, June 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:PBPa

[Ano92i]  Anonymous.  Publications:

Books, periodicals. ACM Fortran Forum, 11(2):6–8, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:PBPb

[Ano92j]  Anonymous.  Publications:

Books, periodicals, software newsletters. ACM Fortran Forum, 11(4):4, December 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:PHP

[Ano92k] Anonymous. Publications: High Performance Fortran, books, periodicals. ACM Fortran Forum, 11(3):5–6, September 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:PRP

[Ano92l] Anonymous. Publications: Reports, periodicals, books. ACM Fortran Forum, 11(1):10–11, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SFSa

[Ano92m] Anonymous. Software: Fortran software. ACM Fortran Forum, 11(1):6–9, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SFSb

[Ano92n] Anonymous. Software: Fortran software. ACM Fortran Forum, 11(2):9–11, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SFSc

[Ano92o] Anonymous. Software: Fortran software. ACM Fortran Forum, 11(3):7–8, September 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SFSd

[Ano92p] Anonymous. Software: Fortran software and Fortran 90 testing and validation and other related software. ACM Fortran Forum, 11(4):5–7, December 1992. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SFSe

[Ano92q] Anonymous. Software: Fortran software; Fortran 90: Fortran 90 testing and validation; Fortran 77; other related software. ACM Fortran Forum, 12(2):6– 8, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIC

[Ano92r] Anonymous. Standards information: Copies of Fortran Standard; interpretations; future meetings; national Fortran working groups; HPFF update; Fortran 90 implementation workshop, Nov 1992. ACM Fortran Forum, 12(2):2–5, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIFd

[Ano92s] Anonymous. Standards information: Fortran 90 text in electronic form; future meetings; national Fortran working groups. ACM Fortran Forum, 11(4):2–3, December 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIFa

[Ano92t] Anonymous. Standards: Information; future meetings; international meetings. ACM Fortran Forum, 11(1):3–4, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIFc

[Ano92u] Anonymous. Standards information: Future meetings; international meetings. ACM Fortran Forum, 11(3):3–4, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1992:SIFb

[Ano92v] Anonymous. Standards: Information; future meetings; international meetings; High Performance Fortran. ACM Fortran Forum, 11(2):2–5, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:CIR

[Ano93a] Anonymous. Correspondence: Input from readers needed. ACM Fortran Forum, 12(3):11, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:FC

[Ano93b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 12(3):9–10, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:PRB

[Ano93c] Anonymous. Publications: Reports, book, periodicals, software newsletters. ACM Fortran Forum, 12(3):10, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:SFS

[Ano93d] Anonymous. Software: Fortran software: Fortran 90; Fortran 90 testing and validation; Fortran 77; other related software. ACM Fortran Forum, 12(3):5– 9, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:HPF

[Ano93e] Anonymous. Special issue: High Performance Fortran language specification, version 1.0 (May 3, 1993) Part I. ACM Fortran Forum, 12(4):i–x, 1–86, December 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). This material also appeared in Scientific Programming, vol. 2, no. 1, June 1993.

Anonymous:1993:SIS

[Ano93f] Anonymous. Standards information: Speak your mind on Fortran 95; interpretations; organization of Standards Committees; future meetings; copies of Fortran 90 Standard; National Fortran working groups; High Performance Fortran language specification. ACM Fortran Forum, 12(3):1–5, September 1993. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:ASC

[Ano94a] Anonymous. ACM SIGPLAN conference proceedings. ACM Fortran Forum, 13(1):31–32, March 1994. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:AJB

[Ano94b] Anonymous. Award for John Backus. ACM Fortran Forum, 13(1):17, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:CIR

[Ano94c] Anonymous. Correspondence: Input from readers needed. ACM Fortran Forum, 13(1):18, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:CRS

[Ano94d] Anonymous. Cray Research shares programming model. ACM Fortran Forum, 13(1):17, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:FCP

[Ano94e] Anonymous. Forthcoming conferences: 1994; 1995; parallel tools consortium organized. ACM Fortran Forum, 13(1):12– 13, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:FCa

[Ano94f] Anonymous. Forthcoming conferences: 1995. ACM Fortran Forum, 13(3):15, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:FCb

[Ano94g] Anonymous. Forthcoming conferences: 1995, 1996. ACM Fortran Forum, 13(4):9, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:PB

[Ano94h] Anonymous. Publications:

Books. ACM Fortran Forum, 13 (3):16, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:PBP

[Ano94i]  Anonymous.  Publications:

Books, periodicals. ACM Fortran Forum, 13(4):10, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:PP

[Ano94j] Anonymous. Publications: Periodicals. ACM Fortran Forum, 13 (3):17, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:PRB

[Ano94k] Anonymous. Publications: Reports; books; periodicals; lecture notes. ACM Fortran Forum, 13 (1):14–15, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SFSc

[Ano94l] Anonymous. Software: Fortran 77 software. ACM Fortran Forum, 13(3):12–14, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SFSb

[Ano94m] Anonymous. Software: Fortran 90 software. ACM Fortran Forum, 13(3):9–11, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SFSa

[Ano94n] Anonymous. Software: Fortran 90 software; Fortran 77 software; other related software. ACM Fortran Forum, 13(1):7– 11, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SFI

[Ano94o] Anonymous. Software: Fortran information resources; Fortran 90 software; Fortran 77 software. ACM Fortran Forum, 13(4):4–8, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIC

[Ano94p] Anonymous. Standards information: Copies of Fortran 90 Standard. ACM Fortran Forum, 13 (3):4, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIFb

[Ano94q] Anonymous. Standards information: Fortran 90 interpretations. ACM Fortran Forum, 13 (3):8, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIFa

[Ano94r] Anonymous. Standards information: Future meetings. ACM Fortran Forum, 13(3):5, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIN

[Ano94s] Anonymous. Standards information: National Fortran working groups. ACM Fortran Forum, 13 (3):5, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIOb

[Ano94t] Anonymous. Standards information: Organization of Standards committees. ACM Fortran Forum, 13(3):4, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIOc

[Ano94u] Anonymous. Standards information: Organization of Standards committees; copies of Fortran 90 Standard; national Fortran working groups; future meetings; plans for Fortran 95; Fortran 90 interpretations. ACM Fortran Forum, 13(4):2–3, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIOa

[Ano94v] Anonymous. Standards information. organization of Standards Committees; Real-Time Fortran Standard; Fortran 90 interpretations; plans for Fortran 95; High Performance Fortran language specification. ACM Fortran Forum, 13(1):4–6, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIPb

[Ano94w] Anonymous. Standards information: Parallel Tools Consortium. ACM Fortran Forum, 13 (3):8, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIPa

[Ano94x] Anonymous. Standards information: Plans for Fortran 95. ACM Fortran Forum, 13(3):6, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIR

[Ano94y] Anonymous. Standards information: Real-time Fortran Standard. ACM Fortran Forum, 13 (3):8, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:ATR

[Ano95a] Anonymous. Announcements: Technical reports to permit Fortran development between formal Standards; HyperLiterature; Lite F90 test suite from Unicomp; ACM SIGForth merged into SIGPlan. ACM Fortran Forum, 14(1):10–11, March 1995.

CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:BRF

[Ano95b] Anonymous. Book reviews: Fortran 90, by Meissner, and Fortran 90 Programming, by Ellis, Philips, and Lahey. ACM Fortran Forum, 14(1):7–9, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FCa

[Ano95c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 14(1):46, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FCb

[Ano95d] Anonymous. Forthcoming conferences. ACM Fortran Forum, 14(1):46, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FCc

[Ano95e] Anonymous. Forthcoming conferences. ACM Fortran Forum, 14(3):21, September 1995. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FCd

[Ano95f] Anonymous. Forthcoming conferences. ACM Fortran Forum, 14(4):20–21, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FR

[Ano95g] Anonymous. Fortran 90 resources. ACM Fortran Forum, 14 (1):42–45, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FSOb

[Ano95h] Anonymous. Fortran Standards: Organization of Standards committees; future meetings. ACM Fortran Forum, 14(4):27, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:FSOa

[Ano95i] Anonymous. Fortran Standards: Organization of Standards committees; future meetings; national Fortran working groups. ACM Fortran Forum, 14(3):35– 36, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:P

[Ano95j] Anonymous. Publications. ACM Fortran Forum, 14(4):19, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:PNF

[Ano95k] Anonymous. Publications: New Fortran books, Fortran 90 classics, other books, periodicals. ACM Fortran Forum, 14(3):22, 24, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:RFF

[Ano95l] Anonymous. REALLOCATE facility for Fortran 2000. ACM Fortran Forum, 14(3):19, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SFR

[Ano95m] Anonymous. Software: F77 resources: F77 compilers; F77 compilers for PC; F77 compilers for Macintosh; F77 code analysis; Toolpack; NAGWare F77 tools; F77 prettyprinters and flowchart generators; other F77 software; FLIB (freeware); UCBTest. ACM Fortran Forum, 14(3):30–34, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SFI

[Ano95n] Anonymous. Software: Fortran information resources; F90 and F95 resources; the F95 Draft Standard; F90 and F95 compilers; F90 tools; other tools; F90 preprocessors; F90 libraries and utilities; F90 tests and benchmarks; Parallel F90. ACM Fortran Forum, 14(3):25– 30, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SS

[Ano95o] Anonymous. Software in the spotlight: Fortran 90 for Windows: Salford FTN90, Lahey LF90 v2.0, Microsoft Fortran Power Station 4.0; Digital Equipment Corp. restructures Fortran product line; plusFort programmer’s toolkit for Linux. ACM Fortran Forum, 14(4):4–6, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SSE

[Ano95p] Anonymous. Software spotlight: ELF 90 — an academic subset Fortran 90 compiler. ACM Fortran Forum, 14(1):5–6, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SSF

[Ano95q] Anonymous. Software spotlights: F77 to F90 converters; GNU F77 (“g77”) – successor to “f2c”? ACM Fortran Forum, 14(3):4– 8, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SI

[Ano95r] Anonymous. Standards information. ACM Fortran Forum, 14 (1):47–48, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:FCa

[Ano96a] Anonymous. Forthcoming conferences. ACM Fortran Forum, 15(1):14, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:FCb

[Ano96b] Anonymous. Forthcoming conferences. ACM Fortran Forum, 15(2):30, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:FSOa

[Ano96c] Anonymous. Fortran Standards:

Organization of Standards committees; future meetings. ACM Fortran Forum, 15(1):20, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:FSOb

[Ano96d] Anonymous. Fortran Standards:

Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 15(2):36, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:FSOc

[Ano96e] Anonymous. Fortran Standards:

Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 15(3):ii, December 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:OFV

[Ano96f] Anonymous. Optimization in Fortran vs. C. ACM Fortran Forum, 15(1):8, April 1996. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:P

[Ano96g] Anonymous. Publications. ACM Fortran Forum, 15(1):13, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:PFC

[Ano96h] Anonymous. Publications: Fortran 90/95 classics; hot off the press: F programming languages textbooks, do-it-yourself ELF textbook, parlez-vous Français?, Numerical Recipes, algorithms; Fortran-oriented periodicals. ACM Fortran Forum, 15(2):27–29, August 1996. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:SFP

[Ano96i] Anonymous. Software: Fortran 90 for Power Macintosh; Fortran 90 and HPF for Linux (PC); Fortran-oriented periodicals; Fortran 90 information (April) from Michael Metcalf. ACM Fortran Forum, 15(1):15– 19, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:SS

[Ano96j] Anonymous. Software in the spotlight: F: What is F?, language features, comparison between F^{T M} and Lahey ELF 90 (Essential Fortran), F summary. ACM Fortran Forum, 15(1):4– 5, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:SSF

[Ano96k] Anonymous. Software in the spotlight: fpp, a new implementation of an old preprocessor; more about F; philosophy of Elf90 — a more disciplined Fortran. ACM Fortran Forum, 15 (2):4–6, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1996:SNA

[Ano96l] Anonymous. Software: New announcements: Free Fortran Subset ELF90; Fortran Plus from N. A. Software; ForeSys tool suite from Scientific Services; Fortran information resources; Fortran 90 information (August) from Michael Metcalf. ACM Fortran Forum, 15(2):31–35, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:AFO

[Ano97a] Anonymous. Announcements:

Fortran 90 for OpenGL. ACM Fortran Forum, 16(2):4, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:AFP

[Ano97b] Anonymous. Announcements:

Fortran 95 published; Fortran Futures 98 scheduled for London in May 1998; Fujitsu and Lahey announce Fortran alliance; progress on C and C++ Standards; Michael Metcalf moving to Bonn; modules: Separating specification from implementation. ACM Fortran Forum, 16 (3):2, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:FC

[Ano97c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 16(3):232–23, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Staff:1997:FS

[Ano97d] Anonymous. Fortran Standards: Final content of Fortran 2000; object model for Fortran 2000: Type extension, objects (polymorphic variables), generic resolution and polymorphic dummy arguments, tagged pointers; organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 16(1):2–6, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:FSOa

[Ano97e] Anonymous. Fortran Standards: Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 16(2):31, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:FSOb

[Ano97f] Anonymous. Fortran Standards: Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 16(3):32, December 1997. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:P

[Ano97g] Anonymous. Publications. ACM Fortran Forum, 16(3):23, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:PFCa

[Ano97h] Anonymous. Publications: Fortran 90 classics; new publications; F and ELF90 books; other books; links to on-line Fortran documents; Fortran K simulates parallel execution; courses; conferences. ACM Fortran Forum, 16(1):24–28, April 1997. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:PFCb

[Ano97i] Anonymous. Publications: Fortran 90 classics; new publications; other books. ACM Fortran Forum, 16(2):23–24, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:SST

[Ano97j] Anonymous. Software in the spotlight: Tools for conversion to Fortran 90 and 95; user experience with DIGITAL VisualFortran 5.0. ACM Fortran Forum, 16(2):2–3, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:SSV

[Ano97k] Anonymous. Software in the spotlight: Visual Fortran for Windows 95 and NT on Intel PC, from Digital Equipment Corp.: Approved by Microsoft for science-engineering migration from PowerStation 4.0; the Fortran processor; Developer Studio. ACM Fortran Forum, 16 (1):7, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:SNPb

[Ano97l] Anonymous. Software: New products; Fortran information resources. ACM Fortran Forum, 16(2):23–30, August 1997. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:SNPc

[Ano97m] Anonymous. Software: New products; Fortran information resources. ACM Fortran Forum, 16(3):24–31, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1997:SNPa

[Ano97n] Anonymous. Software: New products; Fortran information resources; Michael Metcalf’s Fortran 90 information file. ACM Fortran Forum, 16(1):29–34, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:AOO

[Ano98a] Anonymous. Announcement: Object-oriented proposal from French Fortran Standards group. ACM Fortran Forum, 17(1):2, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:ANO

[Ano98b] Anonymous. Announcements:

New official Fortran technical reports; working group 5 documents; OpenGL Fortran 95 bindings; MPI module provides enhanced Fortran support; variable precision arithmetic; Fortran information sites; new Fortran compiler versions from Lahey and Fujitsu; downloadable advanced Fortran textbook; Fortran engineering textbook. ACM Fortran Forum, 17(3):1–2, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:FC

[Ano98c] Anonymous. Forthcoming conferences. ACM Fortran Forum, 17(1):21, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:FFF

[Ano98d] Anonymous. Fortran Futures ’98: Fortran for high performance computing. ACM Fortran Forum, 17(3):4, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:FSOa

[Ano98e] Anonymous. Fortran Standards:

Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 17(1), April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:FSOb

[Ano98f] Anonymous. Fortran Standards: Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 17(2), August 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:FSOc

[Ano98g] Anonymous. Fortran Standards: Organization of Standards committees; future meetings; national Fortran activities. ACM Fortran Forum, 17(3):3, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:P

[Ano98h] Anonymous. Publications. ACM Fortran Forum, 17(1):20–21, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1998:SNP

[Ano98i] Anonymous. Software: New products; Fortran information resources. ACM Fortran Forum, 17(1):22–27, April 1998. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:FAC

[Ano99a] Anonymous. First announcement and call for papers: HPCN Europe 2000. ACM Fortran Forum, 18(3):25, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:FSA

[Ano99b] Anonymous. Fortran Standards activities: Organization of Standards committees; international meetings; the Fortran and HPF Standards. ACM Fortran Forum, 19(1):30–31, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:LFR

[Ano99c] Anonymous. Letter from France [remarks on ISO/WG5 discussions on the future of Fortran]. ACM Fortran Forum, 18(2):1, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:PAa

[Ano99d] Anonymous. Product announcements. ACM Fortran Forum, 18 (2):24–25, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1999:PAb

[Ano99e] Anonymous. Product announcements. ACM Fortran Forum, 18 (3):12, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:ICC

[Ano00a] Anonymous. The 2001 International Conference on Computational Science. ACM Fortran Forum, 19(3):25–26, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:A

[Ano00b] Anonymous. Announcements. ACM Fortran Forum, 19(3):23– 25, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:FCT

[Ano00c] Anonymous. Fortran compiler in the “Top Ten Algorithms”. ACM Fortran Forum, 19(2):1, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:FSOa

[Ano00d] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards. ACM Fortran Forum, 19(2):31– 32, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2000:FSOb

[Ano00e] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards; UK National Activity Report presented to the meeting of ISO/IEC JTC1/SC22/WG5 (Fortran), Oulu, Finland, August 14–18, 2000. ACM Fortran Forum, 19(3):29–31, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:A

[Ano01a] Anonymous. Announcements.

ACM Fortran Forum, 20(3):28, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:AFF

[Ano01b] Anonymous. Announcements:

FCAT (Fortran Coverage Anal-

ysis Tool); coco.f90; f90SQL;

Compaq Visual Fortran Educational Software Donation Program. ACM Fortran Forum, 20 (2):7–8, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:AFP

[Ano01c] Anonymous. Announcements:

Free public beta test of Compaq Fortran 1.1; new release of coco; Lahey/Fujitsu Fortran 95 (LF95) Linux version 6.0; NAG-Ware f95 compiler release 4.1; N. A. Software FortranPlus compiler version 2.2; the Fortran Company; new journal: Journal of Computational Methods in Applied Science and Engineering. ACM Fortran Forum, 20 (1):21–24, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:CR

[Ano01d] Anonymous. Computing reviews. ACM Fortran Forum, 20(3):1, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FFQ

[Ano01e] Anonymous. Fortran Forum questionnaire. ACM Fortran Forum, 20(2):1, August 2001. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSA

[Ano01f] Anonymous. Fortran Standards activities: Organization of Standards committees; international meetings; the Fortran and HPF Standards. ACM Fortran Forum, 20(2):8–9, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSOa

[Ano01g] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards. ACM Fortran Forum, 20(1):31– 32, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2001:FSOb

[Ano01h] Anonymous. Fortran Standards: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF Standards.

ACM Fortran Forum, 20(3):31– 32, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:Aa

[Ano02a] Anonymous. Announcements.

ACM Fortran Forum, 21(1):28– 29, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:Ab

[Ano02b] Anonymous. Announcements.

ACM Fortran Forum, 21(3):24– 26, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:FSAb

[Ano02c] Anonymous. Fortran Standard activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 21 (3):30–31, December 2002. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2002:FSAa

[Ano02d] Anonymous. Fortran Standards activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 21 (1):31–32, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). URL

ftp://softlib.rice.edu/pub/
HPF/hpf-v20.ps.gz;   http:
//www.crpc.rice.edu/HPFF/

                                        ; http://www.fortran.com/
                                        fortran/;   http://www.iso.
                                        ch/; http://www.j3-fortran.
                                        org/;   http://www.nag.co.
                                        uk/sc22wg5/.

Anonymous:2003:A

[Ano03a] Anonymous. Announcements. ACM Fortran Forum, 22(1):22– 24, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2003:FSA

[Ano03b] Anonymous. Fortran Standard activities: Organization of Standards committees; international meetings; national Fortran activities; the Fortran and HPF standards. ACM Fortran Forum, 22 (1):27–28, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2006:FN

[Ano06a] Anonymous. Fortran 2008 news. ACM Fortran Forum, 25(2):1, August 2006. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2006:JNM

[Ano06b] Anonymous. J3 news: Mallory North to retire. ACM Fortran Forum, 25(2):1–2, August 2006. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2006:SNI

[Ano06c] Anonymous. Software news: Intel, PGI, and Zeus. ACM Fortran Forum, 25(2):1, August 2006. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2009:ICL

[Ano09] Anonymous. An interchange on

comp.lang.fortran. ACM For-

tran Forum, 28(3):25–31, December 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:2017:N

[Ano17] Anonymous. News. ACM Fortran Forum, 36(1):1–15, April 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

X3J3:1995:SIF

[ANS95] ANSI Accredited Technical Subcommittee X3J3. Special issue: Fortran 95, committee draft, May 1995 (X3J3/95-007r1, SC33/WG5-N1122). ACM Fortran Forum, 14(2):xviii + 354, June 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Appleyard:2001:CFC

[App01] John Appleyard. Comparing Fortran compilers. ACM Fortran Forum, 20(1):6–10, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Appleyard:2014:FCC

[App14] John Appleyard. Fortran compiler comparisons. ACM Fortran Forum, 33(3):18–32, December 2014. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Appleyard:2015:FCC

[App15] John Appleyard. Fortran compiler comparisons. ACM Fortran Forum, 34(1):31–40, April 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Appleyard:2016:FCC

[App16] John Appleyard. Fortran compiler comparisons 2016–2017. ACM Fortran Forum, 35(3):10– 28, December 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1987:FDF

[AR87] Jeanne Adams and John Reid. Fortran 8X — the draft Fortran standard revision. ACM Fortran Forum, 6(2):27–38, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Armstrong:2012:ISF

[Arm12] Jeffrey B. Armstrong. Introducing simply Fortran. ACM Fortran Forum, 31(2):20–27, August 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1989:SCF

[AW89] Jeanne C. Adams and Jerrold L. Wagener. Summary of changes to Fortran draft. ACM Fortran Forum, 8(3):13–19, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Adams:1982:AAF

[AWP82] Jeanne Adams, Jerrold Wagener, and George Paul Jr. Announcement: ANSI Fortran Technical Committee X3J3 and E G and G Idaho. ACM ForTec Forum, 3 (2):45, June 1982. CODEN ???? ISSN 0735-3731.

Bader:2006:DPF

[Bad06] Reinhold Bader. Dynamic polymorphism in Fortran 95: how submodules can help. ACM Fortran Forum, 25(2):3–7, August 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bader:2007:CDP

[Bad07a] Reinhold Bader. Correction on “Dynamic Polymorphism in Fortran 95” from August 2006 issue. ACM Fortran Forum, 26 (1):34, April 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bader:2007:FBG

[Bad07b] Reinhold Bader. A Fortran binding for the GNU Scientific Library. ACM Fortran Forum, 26 (2):4–11, August 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bader:2013:EIC

[Bad13] Reinhold Bader. Extended interoperation with C. ACM Fortran Forum, 32(3):6–20, December 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bagheri:1992:PFP

[BCS92] Babak Bagheri, Terry W. Clark, and L. Ridgeway Scott. PFortran: a parallel dialect of Fortran. ACM Fortran Forum, 11(3):20–31, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Beebe:2001:FHC

[Bee01] Nelson H. F. Beebe. Fortran in the heavens (contd.). ACM Fortran Forum, 20(1):10, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Beebe:2002:BIF

[Bee02] Nelson H. F. Beebe. Back issues of Fortran Forum. ACM Fortran Forum, 21(1):6–7, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Berns:1985:C

[Ber85a] Gerald M. Berns. Correspondence. ACM Fortran Forum, 4 (4):9–11, December 1985. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Berns:1985:SFR

[Ber85b] Gerald M. Berns. Significant Fortran reliability problems and what Fortran 8X can do about them. ACM Fortran Forum, 4 (1):25–32, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Baudoin:1983:C

[BFP83] C. Baudoin, Mary Fowler, and Mary A. Patterson. Correspondence. ACM ForTec Forum, 2 (2):4–7, June 1983. CODEN ???? ISSN 0735-3731.

Barkley:1997:CNB

[BGvE^{+}97] John Barkley, Roman Grzonka, Axel vom Endt, David Muxworthy, Tim Zeisloft, Michael Metcalf, Van Snyder, Robert Corbett, Lawrie Schonfelder, Dave Bailey, Pierre Hugonnet, Charles Knoebel, Clifford Blair, Loren P. Meissner, Phillip Helbig, Dirk Thorsten Vogel, Ronald Sverdlove, Dick Hendrickson, Malcom Cohen, Roger Young, James Giles, Larry Rolison, Richard Maine, Henry Zongarro, Petros Dafniotis, Jing Guo, David C. P. LaFrance-Linden, Harvey Richardson, Walt Brainerd, Patrice Lignelet, Jean Vezina, David Vowles, Clive Page, Viggo Norum, Neil N. Carlson, John Reid, Mike Lijewski, William F. Mitchell, and Robert Ferrell. Captured on the Net: Brian Meek (obituary); honors for Betty Holberton; significant new applications for Fortran 90 and Fortran 95?; why don’t academic snobs teach Fortran? is object-oriented Fortran a bad idea? status of High Performance Fortran; replacement for data statements; array of pointers; blanks in short form read and print; undefined pointer association status; visibility of implied do variable in array constructor; efficiency of loops; generic linked lists; module variables and the save attribute; modules and libraries; standardize Subset Fortran?; “mathematically equivalent” expressions, “volatile” variables, and optimization; pointers, targets, and allocatable arrays as function results. ACM Fortran Forum, 16 (2):5–17, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bierman:1995:FAI

[Bie95] Keith Bierman. Fortran access to IEEE 754 exceptions. ACM Fortran Forum, 14(3):15– 18, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bej:1990:C

[BJ90] Mark D. Bej and Russell K. Jones. Correspondence. ACM Fortran Forum, 9(1):22–28, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bleikamp:1996:IFP

[Ble96] Richard Bleikamp. Implementation of Fortran 90 pointers. ACM Fortran Forum, 15(1):7, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Blevins:2009:GLL

[Ble09] Jason R. Blevins. A generic linked list implementation in Fortran 95. ACM Fortran Forum, 28(3):2–7, December 2009. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Barton:1982:C

[BPP^{+}82] Bruce J. Barton, Doug Pearson, Ian C. Pyle, Robert C. Groman, and Heinrich Hartl. Correspondence. ACM ForTec Forum, 1 (2):2–8, December 1982. CO-DEN ???? ISSN 0735-3731.

Brainerd:1996:MF

[Bra96] Walt Brainerd. Modules in F.

ACM Fortran Forum, 15(2):14, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Brokate:1992:CRV

[Bro92] K. Brokate. Correspondence:

Re: “Varying Length Strings” (draft, Fortran Forum, June 1992). ACM Fortran Forum, 12(2):12, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Berg:1997:CNF

[BRZ^{+}97] Peer Berg, Bob Runyan, Henry Zongaro, Steve Lionel, Stu Anderson, Alan Miller, Albert Fasso, Phillip Helbig, and Loren Meissner. Captured on the Net: Fortran random number generators, intrinsic and otherwise. ACM Fortran Forum, 16(3):3–4, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Burch:1987:TPF

[Bur87] Carl Burch. Thoughts on a pointer facility. ACM Fortran Forum, 6(2):18–20, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Burnus:2009:FP

[Bur09] Tobias Burnus. Fortran 2003 & 2008 poll. ACM Fortran Forum, 28(2):21–24, August 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Bush:2010:VHE

[Bus10] Ian J. Bush. The view from the high end: Fortran, parallelism and the HECToR service. ACM Fortran Forum, 29(3):3– 10, December 2010. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Busby:2017:NCF

[Bus17] L. E. Busby. A note on compiling Fortran. ACM Fortran Forum, 36(3):4–11, December 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Caffin:1984:MFC

[Caf84] R. N. Caffin. More on Fortran coding conventions. ACM Fortran Forum, 3(3):43–46, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Car:2010:RCI

[Car10] David Car. A reference counting implementation in Fortran 95/ 2003. ACM Fortran Forum, 29 (1):21–28, April 2010. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:1998:CNW

[CBFM^{+}98] Ian Chivers, John Blair-Fish, Steven Morgan, Don Shirer, David Vowles, Shaun Forth, Werner W. Schulz, Adriaan Joubert, Clive Page, Pierre Hugonnet, Walt Brainerd, John Lindsay, Chuckson M. Yokota, Loren P. Meissner, David Galles, Alan Miller, and Steve Lionel. Captured on the Net: Who uses Fortran? the SAVE attribute for local variables; aliasing: Same actual argument corresponds to two dummies. ACM Fortran Forum, 17(1):5–10, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2020:DCC

[Chi20] Ian D. Chivers. The diagnostic capability of the Cray, gfortran, Intel, Nag and Oracle Fortran compilers: Revision 1. ACM Fortran Forum, 39(1):3–7, March 2020. ISSN 1061-7264 (print), 1931-1311 (electronic). URL

https://dl.acm.org/doi/10.
1145/3432987.3432990.

Chmielewksi:2001:MCC

[Chm01a] Karsten Chmielewksi. More on the ‘compilation cascade’. ACM Fortran Forum, 20(2):6, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chmielewski:2001:MCC

[Chm01b] Karsten Chmielewski. More on the ’compilation cascade’. ACM Fortran Forum, 20(2):6– 10, August 2001. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Cisneros:1994:BRM

[Cis94] Gerardo Cisneros. Book review: Migrating to Fortran 90, by James F. Kerrigan. ACM Fortran Forum, 13(1):15–17, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Car:2010:PTC

[CL10] David Car and Michael List. PyF95++: a templating capability for the Fortran 95/2003 language. ACM Fortran Forum, 29 (1):2–20, April 2010. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Clarke:1982:USF

[Cla82] Alan Clarke. Using Standard Fortran — past, present, and future. ACM ForTec Forum, 1(2): 13–15, December 1982. CODEN ???? ISSN 0735-3731.

Clarke:1983:USF

[Cla83] Alan Clarke. Using Standard Fortran — past, present, and future — continue from Vol. 1 No.

2. ACM ForTec Forum, 2(2): 10–15, June 1983. CODEN ???? ISSN 0735-3731.

Clerman:1999:FGV

[Cle99] Norman S. Clerman. Fortran 95 ‘gotcha!’ (variation on a theme). ACM Fortran Forum, 18(3):29– 30, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Clerman:2009:NCA

[Cle09] Norman S. Clerman. Note on creating an array of procedure pointers. ACM Fortran Forum, 28(1):5–12, April 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Clodius:1994:COO

[Clo94] William Clodius. Correspondence: Object-oriented Fortran. ACM Fortran Forum, 13(3):19– 21, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Clodius:1999:FSA

[Clo99] William B. Clodius. Fortran Standard activities: Standard Fortran’s future; organization of Standards committees; international meetings; national Fortran activities. ACM Fortran Forum, 18(2):27–32, August 1999. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Cohen:1997:OOF

[Coh97] Malcolm Cohen. Object orientation and Fortran 2002: Part I. ACM Fortran Forum, 16(3): 5–10, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Cohen:1999:OOF

[Coh99] Malcolm Cohen. Object orientation and Fortran 2002: Part II. ACM Fortran Forum, 18 (1):14–21, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

CBEMA:1989:FD

[Com89] Computer and Business Equipment Manufacturers Association (CBEMA). Fortran 8X draft, May 1989: American National Standard for Information Systems Programming Language Fortran, S8 (X3.9-198X), Revision of X3.9-1978. ACM Fortran Forum, 8(4):various, December 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Cohen:1998:FFA

[CRC^{+}98] Malcolm Cohen, John Reid, Anthony Colyandro, Tim Hopkins, Stuart Purves, Venkat V. S. S. Sastry, Andy Smethurst, Mark Stevens, M. Cross, S. Johnson, C. Ierotheou, P. Leggett, E. Evans, Mishi Derakhshan, Bruce Greer, P. R. Maarleveld, S. W. de Leeuw, H. J. Sips, R. E. H. Kurver, Douglas Miles, V. B. Muchnick Mircea Chis Ster, A. Shafarenko, Bill Walster, M. Bartholomew-Biggs, Christèle Faure, John Pryce, Bob Brode, Mike Hooper, Ian Hounam, and Tony Kehoe. Fortran Futures: Abstracts. ACM Fortran Forum, 17(3):13–20, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2005:FAE

[CS05a] Ian D. Chivers and Jane Sleightholme. Fortran aware editors and development environments. ACM Fortran Forum, 24 (3):12–14, December 2005. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2005:PP

[CS05b] Ian D. Chivers and Jane Sleightholme. Parallel programming. ACM Fortran Forum, 24 (3):15–20, December 2005. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2006:CSD

[CS06a] Ian D. Chivers and Jane Sleightholme. Compiler switches for developing Fortran programs. ACM Fortran Forum, 25(3):5–8, December 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2006:FRC

[CS06b] Ian D. Chivers and Jane Sleightholme. Fortran resources: compilers. ACM Fortran Forum, 25(3):9–13, December 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2007:CSFa

[CS07a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 standard. ACM Fortran Forum, 26(1):7– 9, April 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2007:CSFb

[CS07b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 standard.

ACM Fortran Forum, 26(3):20– 22, December 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2007:INF

[CS07c] Ian D. Chivers and Jane Sleightholme. Introduction to NAG Fortran builder. ACM Fortran Forum, 26(2):12–24, August 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2008:CSFa

[CS08a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 standard. ACM Fortran Forum, 27(2):22– 24, August 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2008:CSFb

[CS08b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 standard. ACM Fortran Forum, 27(3):29– 31, December 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2009:CSFb

[CS09a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards. ACM Fortran Forum, 28(2):15–20, August 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2009:CSF

[CS09b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards. ACM Fortran Forum, 28(3):32–37, December 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2009:CSFa

[CS09c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 standard. ACM Fortran Forum, 28(1):26– 28, April 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2009:ISS

[CS09d] Ian D. Chivers and Jane Sleightholme. An introduction to Sun Studio. ACM Fortran Forum, 28(1):13–25, April 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2010:CSF

[CS10a] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 Standards Revision 6. ACM Fortran Forum, 29 (3):26–34, December 2010. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2010:CSFa

[CS10b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards. ACM Fortran Forum, 29(1):29–34, April 2010.

CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2010:CSFb

[CS10c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards. ACM Fortran Forum, 29(2):28–35, August 2010. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2011:CSFa

[CS11a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 7. ACM Fortran Forum, 30(1):16–24, April 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2011:CSFb

[CS11b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 7. ACM Fortran Forum, 30(2):28–37, August 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2011:CSFc

[CS11c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 8. ACM Fortran Forum, 30(3):10–20, December 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2012:CSFb

[CS12a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 10. ACM Fortran Forum, 31(2):28–39, August 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2012:CSF

[CS12b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 Standards Revision 11. ACM Fortran Forum, 31(3):17–28, December 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2012:CSFa

[CS12c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 9. ACM Fortran Forum, 31(1):23–33, April 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2013:CSFc

[CS13a] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 standards revision 14. ACM Fortran Forum, 32(3): 21–34, December 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2013:CSFa

[CS13b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 12. ACM Fortran Forum, 32(1):8–19, April 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2013:CSFb

[CS13c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 Standards revision 12. ACM Fortran Forum, 32(2):14–26, August 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2014:CSF

[CS14] Ian Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 15. ACM Fortran Forum, 33(2): 38–51, August 2014. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2015:CSFb

[CS15a] Ian Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 standards revision 17. ACM Fortran Forum, 34(3): 7–27, December 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2015:CSFa

[CS15b] Ian D. Chivers and Jane Sleightholm. Compiler support for the Fortran 2003 and 2008 Standards Revision 16. ACM Fortran Forum, 34(2):41–57, August 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2015:INF

[CS15c] Ian D. Chivers and Jane Sleightholme. An introduction to NAG Fortran Builder. ACM Fortran Forum, 34(2):11–40, August 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2016:CSFa

[CS16a] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 standards revision 18. ACM Fortran Forum, 35(1):9–29, April 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2016:CSFb

[CS16b] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 Standards Revision 19. ACM Fortran Forum, 35 (2):28–49, August 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2016:CSFc

[CS16c] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 Standards Revision 20. ACM Fortran Forum, 35 (3):29–50, December 2016. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2017:CSFa

[CS17a] Ian Chivers and Jane Sleightholme.

Compiler support for the Fortran 2003 and 2008 Standards Revision 21. ACM Fortran Forum, 36 (1):21–42, April 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2017:CSFb

[CS17b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 Standards Revision 22. ACM Fortran Forum, 36(3):12–33, December 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2018:CSFb

[CS18a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003, 2008, TS29113, and 2018 Standards Revision 24. ACM Fortran Forum, 37(2):19–41, August 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2018:CSF

[CS18b] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003, 2008, TS29113, and 2018 Standards Revision 25. ACM Fortran Forum, 37(3):8–31, December 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2018:CSFa

[CS18c] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003 and 2008 Standards Revision 22. ACM Fortran Forum, 37(1):44– 67, April 2018. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2019:CSF

[CS19a] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2003, 2008, TS29113, and 2018 Standards Revision 26. ACM Fortran Forum, 38(2):7–36, August 2019. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2019:DCC

[CS19b] Ian D. Chivers and Jane Sleightholme. The diagnostic capability of the Cray, gfortran, Intel, Nag and Oracle Fortran compilers. ACM Fortran Forum, 38(3):4–9, December 2019. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Chivers:2020:CSF

[CS20] Ian D. Chivers and Jane Sleightholme. Compiler support for the Fortran 2008 and 2018 Standards: Version 1. ACM Fortran Forum, 39(1):8–27, March 2020. ISSN 1061-7264 (print), 1931-1311 (electronic). URL

https://dl.acm.org/doi/10.
1145/3432987.3432991.

Curcic:2018:PFF

[Cur18] Milan Curcic. A parallel Fortran framework for neural networks and deep learning. ACM Fortran Forum, 38(1):4–21, December 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

DEntremont:1985:C

[DB85] Edward J. D’Entremont and Kathleen L. Briggs. Correspondence. ACM Fortran Forum, 4 (2):12–15, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Decyk:2008:MPD

[Dec08] Viktor K. Decyk. A method for passing data between C and opaque Fortran 90 pointers. ACM Fortran Forum, 27 (2):2–7, August 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Dedo:1995:NFF

[Ded95] Craig T. Dedo. New features in Fortran 95 (draft of RATIO-NALES appendix). ACM Fortran Forum, 14(1):37–39, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Dedo:1999:DMA

[Ded99] Craig T. Dedo. Debunking the myths about Fortran. ACM Fortran Forum, 18(2):12–21, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Dellinger:1993:CFH

[Del93] Joe Dellinger. Correspondence: The Fortran-hating gateway. ACM Fortran Forum, 12(3):12, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Delves:2000:MUI

[Del00a] L. M. Delves. Making use of IEEE arithmetic facilities. ACM Fortran Forum, 19(3):9–12, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Delves:2000:ITU

[Del00b] Mike Delves. If it’s there, use it. ACM Fortran Forum, 19 (1):13–18, April 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Decyk:1999:FGPb

[DN99] Viktor K. Decyk and Charles D.

Norton. Fortran 90 ‘Gotchas!’ (part 2). ACM Fortran Forum, 18(3):26–28, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Decyk:2000:FGP

[DN00a] Viktor K. Decyk and Charles D.

Norton. Fortran 90 ‘Gotchas!’ (part 3). ACM Fortran Forum, 19(1):19–22, April 2000. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). See correction [DN00b].

Decyk:2000:GC

[DN00b] Viktor K. Decyk and Charles D.

Norton. ‘gotcha’ correction. ACM Fortran Forum, 19(3):8, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). See [DN00a].

Decyk:1997:EOO

[DNS97] Viktor K. Decyk, Charles D. Norton, and Boleslaw K. Szymanski. Expressing objectoriented concepts in Fortran 90. ACM Fortran Forum, 16(1):13– 18, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Decyk:1999:FGPa

[DNS99] Viktor K. Decyk, Charles D. Norton, and Boleslaw K. Szymanski. Fortran 90 ‘Gotchas’ (part 1). ACM Fortran Forum, 18(2):22–23, August 1999. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Dolan:1988:FD

[Dol88] Kent Paul Dolan. Fortran 8X discussion. ACM Fortran Forum, 7 (1):17–27, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2001:CUNa

[Don01a] Aleksandar Donev. Communicating with the user in numerical routines: Part I: OOP approach in F2x. ACM Fortran Forum, 20 (2):11–18, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2001:CUNb

[Don01b] Aleksandar Donev. Communicating with the user in numerical routines: Part II: An iterative linear solver in F95 and F2X. ACM Fortran Forum, 20 (2):19–32, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2002:MNR

[Don02] Aleksandar Donev. Multiple nonzero-rank part references for structure components in Fortran 95/2002. ACM Fortran Forum, 21(3):2–9, December 2002. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2004:GEF

[Don04] Aleksandar Donev. Genericity in extending Fortran. ACM Fortran Forum, 23(1):2–13, April 2004. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2006:ICF

[Don06] Aleksander Donev. Interoperability with C in Fortran 2003. ACM Fortran Forum, 25(1):8– 12, April 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Donev:2007:RCA

[Don07] Aleksandar Donev. Rationale for co-arrays in Fortran 2008. ACM Fortran Forum, 26(3):9– 19, December 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

dePolignac:1996:CN

[dPGS^{+}96] Christian de Polignac, Jing Guo, David Singleton, A. C. Marshall, and John Bray. Captured on the Net. ACM Fortran Forum, 15 (2):2–3, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

DeLong:1992:DEU

[DTF92] Lewis L. DeLong, David B. Thompson, and Janice M. Fulford. Data encapsulation using Fortran 77 modules. ACM Fortran Forum, 11(3):11–19, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Duncan:1993:CES

[DTSB93] Roy Duncan, John Tunstall, Brian T. Smith, and Richard Brankin. Correspondence: Expert systems for re-ordering arithmetic expressions? ACM Fortran Forum, 12(3):12–14, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Dupee:1994:OOM

[Dup94] Brian J. Dupée. Object oriented methods using Fortran 90. ACM Fortran Forum, 13(1):21– 30, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ellis:1996:CNF

[EEV^{+}96] Miles Ellis, Bernd Eggen, Arne Vajhoej, Emili Besalu, and Steve Lionel. Captured on the net: Final Fortran 95 Draft Stan-

dard via ftp; comp-fortran-90

mailbase list now on WWW; HPF home page on WWW; user notes on Fortran programming on WWW; nested summation symbol; session notes for Fall 1995 DECUS DEC Fortran 90 session; ACM groups deplore switch to C++ for advance placement. ACM Fortran Forum, 15 (1):2–3, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Einarsson:2009:SML

[EHH09] Bo Einarsson, Richard J. Hanson, and Tim Hopkins. Standardized mixed language programming for Fortran and C. ACM Fortran Forum, 28(3):8– 22, December 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ellis:1993:CEH

[Ell93] Miles Ellis. Correspondence: Excerpts from the history of Fortran. ACM Fortran Forum, 12 (3):12, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ellis:1998:FFK

[Ell98] Miles Ellis. Fortran futures:

Keynote speakers: Is there a role for Standards in the future of Fortran? ACM Fortran Forum, 17(3):5, December 1998. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Epstein:1996:EMF

[Eps96] David Epstein. Error messages in F. ACM Fortran Forum, 15 (2):15–18, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Epstein:1997:EMF

[Eps97] David Epstein. Error messages in F, take two. ACM Fortran Forum, 16(3):12–16, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Epstein:1999:MTJ

[Eps99] David Epstein. More than just CoCo. ACM Fortran Forum, 18 (3):15–20, December 1999. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Feldman:1990:FCC

[FGMS90] S. I. Feldman, D. M. Gay, M. W. Maimone, and N. L. Schryer. A Fortran to C converter. ACM Fortran Forum, 9 (2):21–22, October 1990. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Feldman:1991:AFF

[FGMS91] S. I. Feldman, D. M. Gay, M. W. Maimone, and N. L. Schryer. Availability of f2c — a Fortran to C converter. ACM Fortran Forum, 10(2):14–15, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Fanfarillo:2015:LOS

[FR15] Alessandro Fanfarillo and Damian Rouson. Leveraging OpenCoarrays to support Coarray Fortran on IBM Power8E. ACM Fortran Forum, 34(2):4–10, August 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Frisbie:1997:CNI

[FRM^{+}97] Ralph Jay Frisbie, Furue Ryo, Loren P. Meissner, Phillip Helbig, William Clodius, Pierre Hugonnet, Kurt Hirchert, Ronald Sverdlove, Clive Page, Richard Maine, Gordon Sande, and Marc Mausch. Captured on the Net: Interest group for F subset email; this train is bound for F2005? side effects and Fortran 95’s pure functions; upper and lower case; Standard Fortran subset?; global allocatable arrays. ACM Fortran Forum, 16(1):8–12, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Forest:2001:FPP

[FS01] Etienne Forest and Frank Schmidt. The “Full Polymorphic Package” (FPP). ACM Fortran Forum, 20(3):12–17, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Fullerton:1987:ADF

[Ful87] James Fullerton. An alternate design for Fortran 8X. ACM Fortran Forum, 6(3):31–40, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Gardner:2006:CAM

[GD06] Henry J. Gardner and Viktor K.

Decyk. Comments on the Arjen Markus article: Design patterns and Fortran. ACM Fortran Forum, 25(2):8–11, August 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Goodger:1999:FF

[GFT^{+}99] Karen Goodger, Alvaro Fernandez, Jean-Yves Tiller, Bill Moffitt, Dan Nagle, Phillip Helbig, John Bray, Charles Augarde, Friedrich Hertweck, Clive Page, Vittorio Caffa, Steve Morgan, Mandeep Singh, Shaun Forth, Jason Abate, Bob Runyan, David B. Serafini, John P. Grimes, Adriaan Joubert, Bertrand Meltz, Pierre Hugonnet, Walt Brainerd, Jack N. Hatfield, and Charles Norton. Future of Fortran. ACM Fortran Forum, 18(1):24– 30, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Giles:1996:ASS

[Gil96] James Giles. Aliasing, syntax, and string handling in Fortran and in C. ACM Fortran Forum, 15(2):12–13, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Giles:2001:EDT

[Gil01] Jim Giles. Enumerated data types. ACM Fortran Forum, 20 (1):11–15, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Goelz:2008:ATF

[Goe08] Tony Goelz. Absoft Tools Fortran Integrated Development Environment. ACM Fortran Forum, 27(1):6–17, April 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Gorelik:2008:SIP

[Gor08] A. M. Gorelik. Statements and intrinsic procedures in consecutive Fortran standards. ACM Fortran Forum, 27(3):16–28, December 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Gorelik:2014:SDTa

[Gor14a] A. M. Gorelik. Statements, data types and intrinsic procedures in the Fortran standards (1966– 2008). ACM Fortran Forum, 33 (2):8–20, August 2014. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Gorelik:2014:SDTb

[Gor14b] A. M. Gorelik. Statements, data types and intrinsic procedures in the Fortran standards (1966– 2008). ACM Fortran Forum, 33 (3):5–17, December 2014. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Greer:1998:FFP

[Gre98] Bruce Greer. Fortran Futures:

Papers: The most important technical library in the world. ACM Fortran Forum, 17(3):25– 32, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Guinier:1987:FPP

[Gui87] Danile Guinier. Fortran programs on a personal computer. ACM Fortran Forum, 6(2):39– 47, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hatfield:2003:CCA

[Hat03] Jack N. Hatfield. CATO (Computer Algorithm for Trajectory Optimization): An implementation of Fortran 95 ojbect-based programming. ACM Fortran Forum, 22(1):2–7, April 2003. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hartl:1987:C

[HBM87] Heinrich Hartl, Gerald M. Burns, and Robert J. Mercer. Correspondence. ACM Fortran Forum, 6(2):23–26, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hendrickson:1990:CEC

[Hen90]  R. A. Hendrickson.  Corre-

spondence: Efficiency considerations for non-contiguous arrays. ACM Fortran Forum, 9(2):17– 18, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hendrikson:2018:FG

[Hen18]  Dick Hendrikson.  Fortran

geezers. ACM Fortran Forum, 37 (2):2–7, August 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Herbert:1994:CNO

[Her94] Wayne V. Herbert. Correspondence: New object-oriented language proposal (grin)?? ACM Fortran Forum, 13(1):18, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hey:1998:NCF

[Hey98] Tony Hey. New challenges for Fortran in the next millennium. ACM Fortran Forum, 17(3):10– 12, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:HPF

[Hig94a] High Performance Fortran Forum. High Performance Fortran language specification (part III), Version 1.0, May 1993. ACM Fortran Forum, 13(3):22– 55, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHa

[Hig94b] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 5. intrinsic and library procedures. ACM Fortran Forum, 13(2):87–138, June 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHb

[Hig94c] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 6. extrinsic procedures. ACM Fortran Forum, 13 (2):139–142, June 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHc

[Hig94d] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 7. storage and sequence association. ACM Fortran Forum, 13(2):143–150, June 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hirchert:1996:OOP

[Hir96]  Kurt Hirchert.  Object ori-

ented programming and Fortran. ACM Fortran Forum, 15(2):23– 26, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Holmes:1995:PF

[Hol95] Neville Holmes. Programming with Fortran. ACM Fortran Forum, 14(1):29–33, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hollenberg:2000:HEJ

[Hol00] Jaap Hollenberg. HPCN Europe 2001 June 25–27, 2001. ACM Fortran Forum, 19(3):26, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hounam:2005:STF

[Hou05] Ian Hounam. Software tools for Fortran. ACM Fortran Forum, 24(1):6–22, April 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hybl:1987:C

[Hyb87] Albert Hybl. Correspondence. ACM Fortran Forum, 6(3):10– 11, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Iles:1985:C

[IMH85] R. M. J. Iles, Michael Metcalf, and David W. Hamaker. Correspondence. ACM Fortran Forum, 4(1):17–24, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

ISO:1996:TRF

[ISO96a] ISO/IEC JTC1/SC22/WG5 — N1231. Technical report for floating-point exception handling. ACM Fortran Forum, 15 (3):1–28, December 1996. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). 7th October 1996.

ISO:1996:EDT

[ISO96b] ISO/IEC JTC1/SC22/WG5/ N1230. Technical report (ISO/ IEC JTC1 PDTR 22.02.01.04): Enhanced derived type facilities: Allocatable array regularization, allocatable arrays as dummy arguments, allocatable array function results, allocatable array components. ACM Fortran Forum, 15(3):29–38, December 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). 10th October 1996.

Iwashita:2000:RHA

[Iwa00] Hidetoshi Iwashita. Report of HUG2000 [4th annual High Performance Fortran user group meeting, October 19–20, 2000, Tokyo, Japan]. ACM Fortran Forum, 19(3):21–22, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Jonson:1995:FOS

[Jon95] Ralph Jonson. Fortran “objects” should be based on generic functions. ACM Fortran Forum, 14 (3):20, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Kaiser:2003:MCL

[Kai03a] Timothy H. Kaiser. A methodology for creating large modules. ACM Fortran Forum, 22(3):11– 24, December 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Kaiser:2003:SCF

[Kai03b] Timothy H. Kaiser. Semiautomatic characterization of Fortran 90 compilers. ACM Fortran Forum, 22(3):2–10, December 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Keirstead:1985:C

[KH85] Friend H. Keirstead, Jr. and Albert Hybl. Correspondence. ACM Fortran Forum, 4(3):17, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Khalil:1991:VFF

[Kha91] M. R. Khalil. VAX-11 Fortran: a flaw in efficient code generation. ACM Fortran Forum, 10 (2):28, July 1991. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Koelbel:1992:OHP

[Koe92] Charles Koelbel. An overview of High Performance Fortran. ACM Fortran Forum, 11(4):9–16, December 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Koelbel:1998:LYW

[Koe98a] Charles Koelbel. The language of the Year 2000: will it really be Fortran. ACM Fortran Forum, 17(3):6–9, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Koelbel:1998:FFP

[Koe98b] Chuck Koelbel. Fortran Futures:

Panel session: the strengths and weaknesses of Fortran 2000. ACM Fortran Forum, 17(3):21, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Kuhn:1998:FFW

[Kuh98] Bob Kuhn. Fortran Futures:

Workshop: OpenMP for parallel Fortran applications. ACM Fortran Forum, 17(3):22, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Lahey:1987:FS

[Lah87] Thomas M. Lahey. The Fortran 8X standard. ACM Fortran Forum, 6(3):27–30, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Lauson:1992:CFT

[Lau92] H. S. (Hank) Lauson. Correspondence: Fortran 90 trivia. ACM Fortran Forum, 11(1):12, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

List:2011:FUT

[LC11a] Michael List and David Car. A Fortran unit-testing framework utilizing templating and the PyF95++ toolset. ACM Fortran Forum, 30(1):3–15, April 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

List:2011:PRC

[LC11b] Michael List and David Car. A polymorphic reference counting implementation in Fortran 2003. ACM Fortran Forum, 30(2):4– 27, August 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Levine:1994:FLE

[Lev94] David Levine. Fortran language extensions. ACM Fortran Forum, 13(4):11–27, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Long:2016:APF

[Lon16] Bill Long. Additional parallel features in Fortran. ACM Fortran Forum, 35(2):16–23, August 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Landman:2000:PLR

[LP00] Joseph Landman and Piotr Piecuch. Parallelization of a legacy research program using OpenMP. ACM Fortran Forum, 19(2):16–23, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Lutowski:1995:OOS

[Lut95] Rick Lutowski. Object-oriented software development with traditional languages. ACM Fortran Forum, 14(4):13–18, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Maine:1996:AAS

[Mai96] Richard Maine. Allocatable arrays as structure components. ACM Fortran Forum, 15(1):6, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Martin:1992:FPV

[Mar92] Jeanne Martin. Fortran 90 pointers vs. “Cray” pointers. ACM Fortran Forum, 11(2):17– 23, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2001:GPF

[Mar01] Arjen Markus. Generic programming in Fortran 90. ACM Fortran Forum, 20(3):20–23, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2002:CFS

[Mar02] Arjen Markus. Combining Fortran and scripting languages. ACM Fortran Forum, 21(3):10– 18, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2003:AML

[Mar03a] Arjen Markus. Avoiding memory leaks with derived types. ACM Fortran Forum, 22(2):1– 6, August 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2003:RSO

[Mar03b] Arjen Markus. Review of “Object-oriented Programming via Fortran 90/95 by Ed Akin.” Cambridge University Press, 2003. ACM Fortran Forum, 22 (2):7–8, August 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Martin:2003:FTQ

[Mar03c] Jeanne T. Martin. Fortran trivia questions. ACM Fortran Forum, 22(2):18–25, August 2003. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2004:LXF

[Mar04a] Arjen Markus. Have a look at XML files. ACM Fortran Forum, 23(3):2–10, December 2004. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2004:QCF

[Mar04b] Arjen Markus. Quantum computing in Fortran. ACM Fortran Forum, 23(2):5–11, August 2004. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2005:SLX

[Mar05] Arjen Markus. Have a second look at XML files. ACM Fortran Forum, 24(1):2–5, April 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2006:DPF

[Mar06] Arjen Markus. Design patterns and Fortran 90/95. ACM Fortran Forum, 25(1):13–29, April 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2008:DPF

[Mar08a] Arjen Markus. Design patterns and Fortran 2003. ACM Fortran Forum, 27(3):2–15, December 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2008:NAV

[Mar08b] Arjen Markus. Note on arrayvalued functions. ACM Fortran Forum, 27(1):2–5, April 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2009:CR

[Mar09] Arjen Markus. Code reviews.

ACM Fortran Forum, 28(2):4– 14, August 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2013:EHF

[Mar13] Arjen Markus. Exception handling in Fortran. ACM Fortran Forum, 32(2):7–13, August 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2014:RSS

[Mar14a] Arjen Markus. Review: Scientific Software Design by Damian Rouson, Jim Xia and Xiaofeng Xu. ACM Fortran Forum, 33 (1):4–5, April 2014. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2014:UDO

[Mar14b] Arjen Markus. User-defined and overloaded operations. ACM Fortran Forum, 33(1):6–15, April 2014. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2017:PHI

[Mar17] Arjen Markus. Pointers for handling input and output. ACM Fortran Forum, 36(1):16– 20, April 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Markus:2019:EGP

[Mar19] Arjen Markus. Experimenting with generic programming features. ACM Fortran Forum, 38 (3):10–19, December 2019. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1996:FOO

[MCC96] Loren P. Meissner, John S. Cuthbertson, and William B. Clodius. Fortran object oriented programming examples. ACM Fortran Forum, 15(1):11–12, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

McCormack:2005:GPF

[McC05] Drew McCormack. Generic programming in Fortran with Forpedo. ACM Fortran Forum, 24 (2):18–29, August 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Moss:1995:CNF

[MCL^{+}95] Leonard J. Moss, William B.

Clodius, Dave Lloyd, Michael Metcalf, Bernd R. Eggen, Chuck Ritz, Cliver Page, Marty Cohen, Emilio C. Lopes, and Jerrold L. Wagener. Captured on the Net: Fortran is alive and well and living in London; where to find Fortran 90 varying string module; formal grammars for Fortran 77 and Fortran 90; linking modules without MAKE; Fortran 90 efficient storage for triangular matrices; email group welcomes participants; Fortran Journal wants more articles; downloadable Fortran 77 textbook (free); pass values, not variables (objects) as arguments; ELF linker — not to be confused with Essential Lahey Fortran; High Performance Fortran Forum meeting; latest Fortran 95 draft available on line. ACM Fortran Forum, 14(4):1–3, December 1995. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

Murali:1993:IFQ

[MDB93] P. Murali, D. Dutta, and R. N. Biswas. An integrated framework for quality scientific software development. ACM Fortran Forum, 12(3):19–25, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meek:1990:FFS

[Mee90] Brian Meek. The Fortran (not the foresight) Saga: the light and the dark. ACM Fortran Forum, 9(2):23–32, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1982:ENT

[Mei82a] Loren P. Meissner. Editor’s note: Times are changing. ACM ForTec Forum, 1(1):2, July 1982. CODEN ???? ISSN 0735-3731.

Meissner:1982:FPL

[Mei82b] Loren P. Meissner. The Fortran programming language — recent developments and a view of the future. ACM ForTec Forum, 1 (1):3–8, July 1982. CODEN ???? ISSN 0735-3731.

Meissner:1983:EIF

[Mei83a] Loren P. Meissner. From the Editor: Interpretations of Fortran 77. ACM ForTec Forum, 2(3):1, September 1983. CODEN ???? ISSN 0735-3731.

Meissner:1983:ESP

[Mei83b] Loren P. Meissner. From the Editor: Special promotional distribution; questions concerning the Fortran Standard. ACM ForTec Forum, 2(1):1, March 1983. CO-DEN ???? ISSN 0735-3731.

Meissner:1983:EWF

[Mei83c] Loren P. Meissner. From the Editor: Whither ForTec? ACM ForTec Forum, 2(4):1, December 1983. CODEN ???? ISSN 0735-3731.

Meissner:1984:EFF

[Mei84a] Loren P. Meissner. From the Editor: ForTec Forum becomes Fortran Forum; Fortran 77 on micros. ACM Fortran Forum, 3 (3):1, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1984:EVA

[Mei84b] Loren P. Meissner. From the Editor: Vote on ACM constitutional amendment; whither ForTec?; Fortran 77 on micros?; UNIX Fortran 77 rumor. ACM ForTec Forum, 3(1):1–2, March 1984. CODEN ???? ISSN 0735-3731.

Meissner:1985:Ea

[Mei85a] Loren P. Meissner. From the Editor. ACM Fortran Forum, 4(1):1, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1985:Eb

[Mei85b] Loren P. Meissner. From the Editor. ACM Fortran Forum, 4 (2):1, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1985:Ec

[Mei85c] Loren P. Meissner. From the Editor. ACM Fortran Forum, 4(3): 1, 11–12, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1985:FRE

[Mei85d] Loren P. Meissner. On “Fortran reliability” (editorial). ACM Fortran Forum, 4(4):1, December 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1986:FTA

[Mei86] Loren P. Meissner. Fast turnaround (editorial). ACM Fortran Forum, 5(2):1, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1987:GFV

[Mei87a] Loren P. Meissner. The glorious future vs. the good old days (editorial). ACM Fortran Forum, 6(1):1, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1987:SE

[Mei87b] Loren P. Meissner. Speak up! (editorial). ACM Fortran Forum, 6(2):2–3, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1987:CFD

[Mei87c] Lorer P. Meissner. Concerning the Fortran 8X draft (editorial). ACM Fortran Forum, 6 (3):1, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1988:MFP

[Mei88a] Loren Meissner. More on Fortran 8X pointer proposals. ACM Fortran Forum, 7(1):28–36, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1988:WDW

[Mei88b] Loren P. Meissner. Where do we go from here? (editorial). ACM Fortran Forum, 7(1):1, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:ESFb

[Mei89a] Loren P. Meissner. Editorial:

Second Fortran 88 public review announced; Fortran 88 implementation conference. ACM Fortran Forum, 8(3):1–2, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:ESFa

[Mei89b] Loren P. Meissner. Editorial:

Second Fortran 88 public review in Fall 1989 likely; ISO Working Group 55 role increasingly important. ACM Fortran Forum, 8(2):1, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:PRI

[Mei89c] Loren P. Meissner. Partial-record input and output. ACM Fortran Forum, 8(3):20–23, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:SF

[Mei89d] Loren P. Meissner. Summary of Fortran 88. ACM Fortran Forum, 8(2):15–31, May 1989. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1989:XAD

[Mei89e] Loren P. Meissner. X3J3 avoids deadlock. ACM Fortran Forum, 8(1):1–3, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1990:CCI

[Mei90a] L. P. Meissner. Correspondence: Current implementations of Fortran 90 features. ACM Fortran Forum, 9(2):13–17, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1990:NFS

[Mei90b] Loren Meissner. A note on Fortran 90 source form. ACM Fortran Forum, 9(2):19–20, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1990:ETU

[Mei90c] Loren P. Meissner. From the Editor: Third US public review for Fortran 90; “milestones of a revolution”. ACM Fortran Forum, 9(2):1, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1990:EWG

[Mei90d] Loren P. Meissner. From the Editor: what is going here? 1. distribution of the Fortran draft;

2. naming the baby; 3. double standards; 4. finishing touches; 5. thank you, John. ACM Fortran Forum, 9(1):1–3, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1990:SRF

[Mei90e] Loren P. Meissner. Summary of revised Fortran – Part 2. ACM Fortran Forum, 9(1):31– 39, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1991:EFT

[Mei91a] Loren P. Meissner. From the editor: Fortran 90 is technically complete and red faces department and first Fortran 90 compiler? input from readers needed. ACM Fortran Forum, 10 (2):1, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1991:EWW

[Mei91b] Loren P. Meissner. From the Editor: When will it be over? Fortran 90 implementation; “milestones of a revolution”. ACM Fortran Forum, 10(1):1, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1992:EAA

[Mei92a] Loren P. Meissner. From the Editor: Amazing ANSI approval announcement (ho-hum …); short count; object-oriented Fortran?; new Fortran 90 software. ACM Fortran Forum, 11(4):1, December 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1992:EFSb

[Mei92b] Loren P. Meissner. From the Editor: Fortran 90 Standard and Fortran 90 compilers and forthcoming in Fortran Forum. ACM Fortran Forum, 11(2):1, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1992:EFSa

[Mei92c] Loren P. Meissner. From the Editor: Fortran 90 Standard; Fortran 90 compilers; Department of Defense group proposes rejection of Fortran 90; forthcoming in Fortran Forum; new Fortran Forum subscription rates; input from readers needed. ACM Fortran Forum, 11(1):1–2, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1992:ETD

[Mei92d] Loren P. Meissner. From the Editor: Thoughts for the day; Fortran 90 Standard; Fortran 90 compilers. ACM Fortran Forum, 11(3):1–2, September 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1992:EUF

[Mei92e] Loren P. Meissner. From the Editor: Update on Fortran Forum special issue, March 1993. ACM Fortran Forum, 12(2):1, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1993:EAF

[Mei93] Loren P. Meissner. From the editor: All Fortran 90 lexers should generate optional free source form output. ACM Fortran Forum, 12(3):1, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1994:EFF

[Mei94a] Loren P. Meissner. From the Editor: Fortran Forum schedule change; HPFF 1995 meeting. ACM Fortran Forum, 13 (4):1, December 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1994:EFD

[Mei94b] Loren P. Meissner. From the Editor: Fortran’s dinosaur image. ACM Fortran Forum, 13 (1):1–3, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1994:EJF

[Mei94c] Loren P. Meissner. From the Editor: Jurassic font? Fortran 95 draft is on schedule; Fortran information resources. ACM Fortran Forum, 13(3):1, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1995:EWP

[Mei95a] L. P. Meissner. From the Editor: Who pays for Standards? Who should pay for Standards? ACM Fortran Forum, 14(1):1– 4, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1995:EAD

[Mei95b] Loren P. Meissner. From the Editor: Allocatable dummy argument arrays; how should Fortran Standards describe arithmetic? are external procedures obsolete?; when are local variables initialized in F77 and F90? ACM Fortran Forum, 14(3):1–3, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1996:FCC

[Mei96a] Loren Meissner. Fortran conditional compilation: preliminary specification. ACM Fortran Forum, 15(2):8–11, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1996:EWS

[Mei96b] Loren P. Meissner. Editorial: WG5 sets Fortran 2000 requirements; has OOF missed the F2000 train?; close vote on Fortran conditional compilation facility. ACM Fortran Forum, 15 (2):1–2, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1996:EFF

[Mei96c] Loren P. Meissner. From the Editor: Fortran Forum subscriptions for non ACM members. ACM Fortran Forum, 15(1):1, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1996:SID

[Mei96d] Loren P. Meissner. Special issue:

Draft technical reports: Introduction. ACM Fortran Forum, 15(3):i, December 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1997:LGN

[Mei97a] Loren Meissner. Little giants — the new Fortran subsets. ACM Fortran Forum, 16(1):20– 23, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1997:SLS

[Mei97b] Loren Meissner. Solving linear systems with Fortran 90 and 95. ACM Fortran Forum, 16(3):17– 22, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1997:FLL

[Mei97c] Loren P. Meissner. Fortran 90 & 95 linked list operations: managing an ordered list with pointers to pointers. ACM Fortran Forum, 16(2):18–22, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1998:ECN

[Mei98a] Loren P. Meissner. Editorial: Call for nominations: Forum Editor; Fortran newsletter history. ACM Fortran Forum, 17(1):1– 2, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1998:NWC

[Mei98b] Loren P. Meissner. Newsletter of the West Coast Working Group on Fortran language development. ACM Fortran Forum, 17(1):3, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). Reprint of issue 1 (1 Febrary 1975) of the Fortran newsletter, later called =

F O R - W O R D - >.

Meissner:1999:ETG

[Mei99] Loren P. Meissner. Editorial: Thin on the ground. ACM Fortran Forum, 18(1):1, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1983:FCC

[Met83a] Michael Metcalf. Fortran 77 coding conventions. ACM ForTec Forum, 2(4):10–15, December 1983. CODEN ???? ISSN 0735-

Metcalf:1983:WF

[Met83b] Michael Metcalf. Why Fortran?

ACM ForTec Forum, 2(1):13–14, March 1983. CODEN ???? ISSN 0735-3731.

Metcalf:1984:FF

[Met84] Michael Metcalf. Has Fortran a future? ACM Fortran Forum, 3(3):21–38, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1987:FES

[Met87] Michael Metcalf. Fortran 8X — the emerging standard. ACM Fortran Forum, 6(1):28–47, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1991:CSF

[Met91] Michael Metcalf. Correspondence: Source form conversion. ACM Fortran Forum, 10(1):14– 15, March 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1992:FEF

[Met92] Michael Metcalf. A first encounter with f90. ACM Fortran Forum, 11(1):24–32, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1994:CNR

[Met94] Michael Metcalf. Correspondence: Numerical Recipes. ACM Fortran Forum, 13(3):21, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1995:FIR

[Met95] Michael Metcalf. Fortran information resources. ACM Fortran Forum, 14(4):22–26, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1996:F

[Met96] Michael Metcalf. Fortran 95. ACM Fortran Forum, 15(2):19– 22, August 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1998:FFP

[Met98] Michael Metcalf. Fortran Futures: Papers: The F programming language. ACM Fortran Forum, 17(3):23–24, December 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:Ea

[Met99a] Michael Metcalf. Editorial. ACM Fortran Forum, 18(3):1, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:Eb

[Met99b] Michael Metcalf. Editorial. ACM Fortran Forum, 19(1):1, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:FHIa

[Met99c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 18(1):22–23, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:FHIb

[Met99d] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 2, utilities). ACM Fortran Forum, 18(2):26–27, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:FHIc

[Met99e] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 3, books in English). ACM Fortran Forum, 18(3):31, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:LFP

[Met99f] Michael Metcalf. Letter from France: a postscript. ACM Fortran Forum, 18(3):1, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:NCA

[Met99g] Michael Metcalf. Notice to contributing authors to SIG Newsletters. ACM Fortran Forum, 19(1):1, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1999:NPA

[Met99h] Michael Metcalf. Notice to past authors of ACM-published articles. ACM Fortran Forum, 19 (1):1, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIb

[Met00a] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 19(2):29–30, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIc

[Met00b] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 2, tools). ACM Fortran Forum, 19 (3):27–28, December 2000. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FHIa

[Met00c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 4, courses etc.). ACM Fortran Forum, 19(1):28–29, April 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2000:FFQ

[Met00d] Michael Metcalf. Fortran Forum questionnaire. ACM Fortran Forum, 19(3):1, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:E

[Met01a] Michael Metcalf. Editorial. ACM Fortran Forum, 20(1):1, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIc

[Met01b] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 1, compilers). ACM Fortran Forum, 20(3):29–30, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIa

[Met01c] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 3, books). ACM Fortran Forum, 20 (1):27–30, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FHIb

[Met01d] Michael Metcalf. Fortran 90/ 95/HPF information file (Part 4, courses, etc.). ACM Fortran Forum, 20(2):10, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2001:FFQ

[Met01e] Michael Metcalf. Fortran Forum questionnaire. ACM Fortran Forum, 20(1):1, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:E

[Met02a] Michael Metcalf. Editorial. ACM Fortran Forum, 21(3):1, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:FHIa

[Met02b] Michael Metcalf. Fortran 90/ 95/HPF information file (part 2, tools, etc.). ACM Fortran Forum, 21(1):30–31, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:FHIb

[Met02c] Michael Metcalf. Fortran 90/ 95/HPF information file (part 3, books). ACM Fortran Forum, 21 (3):27–29, December 2002. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:HF

[Met02d] Michael Metcalf. History of Fortran. ACM Fortran Forum, 21 (3):19–20, December 2002. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2002:PPI

[Met02e] Michael Metcalf. Pretty printing and indenting tools. ACM Fortran Forum, 21(1):25–27, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2003:EL

[Met03a] Michael Metcalf. Editorial: On legitimacy. ACM Fortran Forum, 22(1):1, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2003:FIF

[Met03b] Michael Metcalf. Fortran information file (part 4, courses, consultancy, standards). ACM Fortran Forum, 22(1):25–26, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2006:SPF

[Met06] Michael Metcalf. A Sudoku program in Fortran 95. ACM Fortran Forum, 25(1):4–7, April 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2019:HF

[Met19] Michael Metcalf. History of Fortran. ACM Fortran Forum, 38 (2):4–6, August 2019. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:2020:BRH

[Met20] Michael Metcalf. Book review:

The History of the Fortran Programming Language: Abstracting away the machine. ACM Fortran Forum, 39(1):1–2, March 2020. ISSN 1061-7264 (print), 1931-1311 (electronic). URL

https://dl.acm.org/doi/10.
1145/3432987.3432989.

Moniot:1992:C

[MF92] Robert Moniot and Lloyd Fosdick. Correspondence. ACM Fortran Forum, 11(2):15–16, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Mitchell:1999:FBO

[Mit99] William F. Mitchell. The Fortran 90 bindings for OpenGL. ACM Fortran Forum, 18(1):5– 13, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Miyawaki:1987:FA

[Miy87] Stephen M. Miyawaki. Fortran 8X abridgement: ANSI X3J3/S8 Version 104 Working Document, June 1987. ACM Fortran Forum, 6(3):41–78, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Meissner:1984:C

[MLW84] Loren P. Meissner, K. W. Loach, and N. A. Wagg. Correspondence. ACM Fortran Forum, 3 (3):16–20, December 1984. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Moss:1990:LMR

[Mos90] Len Moss. Len Moss reports. ACM Fortran Forum, 9(1):14, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Metcalf:1991:WF

[MR91] Michael Metcalf and John Reid. Whither Fortran? ACM Fortran Forum, 10(2):18–22, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Murphy:1989:C

[Mur89] Harry M. Murphy. Correspondence. ACM Fortran Forum, 8 (2):13–14, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Muxworthy:2009:WMN

[Mux09] David Muxworthy. WG5 meetings: November 2008 and May 2009: a personal view. ACM Fortran Forum, 28(2):2–3, August 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Muxworthy:2013:RJW

[Mux13] David Muxworthy. A report on the joint WG5 and PL22.3 meeting June 2013, Delft, Netherlands. ACM Fortran Forum, 32 (3):4–5, December 2013. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

McGavin:2001:GLI

[MY01] Peter McGavin and Roger Young. A generic list implementation. ACM Fortran Forum, 20 (1):16–20, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

NAG:1985:TFS

[NAG85] NAG. TOOLPACK/1 Fortran software tools suite — summary of contents (February 1985). ACM Fortran Forum, 4(1):39– 40, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:1999:PP

[Nag99] Dan Nagle. The portability project. ACM Fortran Forum, 18 (2):4–7, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2001:IC

[Nag01] Dan Nagle. An implementation of coco. ACM Fortran Forum, 20(1):2–5, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2002:WF

[Nag02] Dan Nagle. Whither Fortran? ACM Fortran Forum, 21(1):1– 5, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2003:WNF

[Nag03] Dan Nagle. What’s new in Fortran 2000? ACM Fortran Forum, 22(2):9–12, August 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2005:AMB

[Nag05] Dan Nagle. Archaic to modern 1: basic tools & issues. ACM Fortran Forum, 24(3):4–7, December 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2006:AMK

[Nag06] Dan Nagle. Archaic to modern 2: kinds and modules. ACM Fortran Forum, 25(3):2–4, December 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nagle:2009:JJW

[Nag09] Dan Nagle. Joint J3/WG5 meetings November 2008, May 2009: a reply to David Muxworthy. ACM Fortran Forum, 28(3):23– 24, December 2009. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Nguyen:2003:SFP

[Ngu03] Thi Viet Nga Nguyen. Sequencing Fortran program array verifications. ACM Fortran Forum, 22(1):8–10, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Numrich:1998:SRC

[NR98] Robert W. Numrich and John Reid. Special report: Co-Array Fortran for parallel programming. ACM Fortran Forum, 17 (2):1–31, August 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Numrich:2005:CAN

[NR05] Robert W. Numrich and John Reid. Co-arrays in the next Fortran Standard. ACM Fortran Forum, 24(2):4–17, August 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

OGara:1995:SET

[OAB^{+}95] Linda O’Gara, Jeanne Adams, Walt Brainerd, Vic Kelson, Craig Dedo, Keith Bierman, Jerry Wagener, Richard Maine, Leonard J. Moss, James H. Billen, Robert Corbett, Dick Hendrickson, Jamie Shiers, David Levine, David L. Epstein, John Reid, Lawrie Schonfelder, Kurt W. Hirchert, Jens Helmers, William Behrman, A. C. Marshall, and William B. Clodius. Some email threads: Is Fortran 90 succeeding?; Fortran market acceptance; Fortran–C interoperability; Fortran preprocessors or conditional compilation: Is standardization needed?; floating point subsets of enable; pointers to procedures, or procedure names as variables; standard linear algebra and advanced math functions modules; re: Allocatable arrays in structures; restricted module visibility. ACM Fortran Forum, 14 (1):11–28, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

OGara:1994:CD

[OCMB94] Linda O’Gara, Roger Caffin, Neldon H. Marshall, and Keith Bierman. Correspondence: Dinosaurs? ACM Fortran Forum, 13(3):18–19, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Overbey:2010:CRS

[OFKJ10] Jeffrey L. Overbey, Matthew J. Fotzler, Ashley J. Kasza, and Ralph E. Johnson. A collection of refactoring specifications for Fortran 95. ACM Fortran Forum, 29 (3):11–25, December 2010. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Olagnon:2001:XXX

[Ola01a] Michel Olagnon. 1000!. ACM Fortran Forum, 20(2):5, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Olagnon:2001:OTF

[Ola01b] Michel Olagnon. 1000! ACM Fortran Forum, 20(2):5, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Page:1996:HSF

[Pag96] Clive Page. How to structure a F90 procedure library. ACM Fortran Forum, 15(1):9–10, April 1996. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Page:2005:SIO

[Pag05] Clive Page. Stream input/output in Fortran. ACM Fortran Forum, 24(3):8–11, December 2005. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Patton:1993:OOF

[Pat93] Bob Patton. Object-oriented Fortran 77 (a practitioner’s view). ACM Fortran Forum, 12 (2):23–24, June 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Payne:1986:C

[PBF86] Fred R. Payne, Gerald M. Berns, and James J. Fullerton. Correspondence. ACM Fortran Forum, 5(1):14–19, July 1986. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Poppe:2012:EHF

[PCV12] Koen Poppe, Ronald Cools, and Bart Vandewoestyne. Error handling in Fortran 2003. ACM Fortran Forum, 31(2):7–19, August 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Software:2012:FCC

[Pol12] Polyhedron Software. Fortran compiler comparisons diagnostic capabilities. ACM Fortran Forum, 31(3):7–16, December 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Polyhedron:2013:FCC

[Pol13] Polyhedron Software. Fortran compiler comparisons execution time benchmarks. ACM Fortran Forum, 32(1):4–7, April 2013. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Prentice:1995:FPS

[Pre95] John K. Prentice. Fortran 90 for physics students. ACM Fortran Forum, 14(1):34–36, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Prince:1999:FPP

[Pri99] Timothy C. Prince. Fortran 90 performance problems. ACM Fortran Forum, 18(3):21–25, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ragan:1982:PAA

[Rag82] Richard R. Ragan. Proposed argument association rules for arrays in Fortran 8X. ACM ForTec Forum, 1(1):9–14, July 1982. CODEN ???? ISSN 0735-3731.

Redish:1991:CBW

[Red91] K. A. Redish. A comparison between Watfor-77 and Fortran 90. ACM Fortran Forum, 10 (2):23–27, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Redwine:1993:DWR

[Red93]  Cooper Redwine.  The DO

WHILE reconsidered. ACM Fortran Forum, 12(3):15–16, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Redwine:1994:DAF

[Red94] Cooper Redwine. Degenerate arrays in Fortran 90. ACM Fortran Forum, 13(1):19–20, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1986:JRRb

[Rei86a] John Reid. John Reid reports:

X3J3 meeting, April 1986. ACM Fortran Forum, 5(1):7–10, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1986:JRRa

[Rei86b] John Reid. John Reid reports: X3J3 meeting, January 1986. ACM Fortran Forum, 5 (1):4–6, July 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1986:JRRc

[Rei86c] John Reid. John Reid reports:

X3J3 meeting, January 1986. ACM Fortran Forum, 5(2):4–8, August 1986. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRa

[Rei87a] John Reid. John Reid reports: X3J3 meeting, August 1986. ACM Fortran Forum, 6 (1):4–7, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRf

[Rei87b] John Reid. John Reid reports: X3J3 meeting, August 1987 (Liverpool). ACM Fortran Forum, 6(2):12–17, October 1987. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRc

[Rei87c] John Reid. John Reid reports: X3J3 meeting, February 1987. ACM Fortran Forum, 6(1):12– 13, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRd

[Rei87d] John Reid. John Reid reports: X3J3 meeting, March 1987 (Albuquerque). ACM Fortran Forum, 6(2):6–7, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRe

[Rei87e] John Reid. John Reid reports: X3J3 meeting, May 1987 (Seattle). ACM Fortran Forum, 6(2): 8–11, October 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:JRRb

[Rei87f] John Reid. John Reid reports:

X3J3 meeting, November 1986. ACM Fortran Forum, 6(1):8– 11, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1987:XMN

[Rei87g]  John Reid.  X3J3 meeting,

November 1987 (Fort Lauderdale). ACM Fortran Forum, 6(3):4–6, December 1987. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1988:JRR

[Rei88] John Reid. John Reid reports: X3J3 meeting, February 1988 (New Orleans). ACM Fortran Forum, 7(1):4–5, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:ERC

[Rei89a] Edward Reid. Edward Reid:

some comments on Fortran 8X. ACM Fortran Forum, 8(1):33– 36, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:ERS

[Rei89b] Edward Reid. Some comments on Fortran 8X. ACM Fortran Forum, 8(1):33–36, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:JRRa

[Rei89c] John Reid. John Reid reports:

X3J3 Meeting 108, 9–13 May 1988, Urbana IL; X3J3 Meeting 109, 8–12 Aug 1988, Jackson WY; ISO/WG5 Meeting, 19– 23 Sep 1988, Paris; X3J3 Meeting 110, 14–18 Nov 1988, Boston MA. ACM Fortran Forum, 8 (1):6–17, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:JRRb

[Rei89d] John Reid. John Reid reports: X3J3 meeting 111, 12–17 Feb 1989, Palo Alto CA. ACM Fortran Forum, 8(2):4–8, May 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1989:JRRc

[Rei89e] John Reid. John Reid reports: X3J3 meeting 112, 8–12 May 1989, Port Jefferson NY. ACM Fortran Forum, 8(3):5–7, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1990:JRR

[Rei90] John Reid. John Reid reports: WG5 meeting, 10–13 Jul 1989, Ispra Italy; X3J3 meeting 113 17–21 July 1989, Vienna Austria; X3J3 meeting, 16–20 Oct 1989, Estes Park, CO; X3J3 meeting 114, 8–12 Jan 1990, Richardson TX. ACM Fortran Forum, 9 (1):6–13, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1992:PPF

[Rei92] John Reid. On PCF parallel Fortran extensions. ACM Fortran Forum, 11(1):17–23, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1993:UFD

[Rei93a] Edward Reid. Using the Fortran 90 DO WHILE. ACM Fortran Forum, 12(2):13–20, June 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1993:DWV

[Rei93b] John Reid. On DO WHILE vs.

DO with EXIT. ACM Fortran Forum, 12(3):17–18, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1995:EHF

[Rei95] John Reid. Exception handling in Fortran. ACM Fortran Forum, 14(3):9–15, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1999:MCW

[Rei99] John Reid. Message from the convenor of WG5. ACM Fortran Forum, 19(1):2, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2000:CF

[Rei00a] John Reid. Content of Fortran 2000. ACM Fortran Forum, 19 (1):3–4, April 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2000:CIF

[Rei00b] John Reid. From the convenor of the International Fortran Standards Committee (WG5): WG5 meeting in Oulu, August 14– 18, 2000; SC22 meeting in Nara, Japan, September 12–15, 2000. ACM Fortran Forum, 19(3):17– 20, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2001:IFS

[Rei01] John Reid. International Fortran standards committee (WG5). ACM Fortran Forum, 20(3):24– 27, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2002:FF

[Rei02a] John Reid. Fortran 2000 forum. ACM Fortran Forum, 21 (3):23, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2002:CIF

[Rei02b] John Reid. From the Convenor of the International Fortran Standards Committee (WG5). ACM Fortran Forum, 21(3):21–22, December 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2002:NFF

[Rei02c] John Reid. The new features of Fortran 2000. ACM Fortran Forum, 21(2):1–31, August 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2003:FCB

[Rei03] John Reid. Fortran 2000 CD ballot and WG5’s response. ACM Fortran Forum, 22(2):13–17, August 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2006:MFN

[Rei06] John Reid. The major features of the next revision of Fortran. ACM Fortran Forum, 25(2):12– 15, August 2006. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2007:NFF

[Rei07] John Reid. The new features of Fortran 2003. ACM Fortran Forum, 26(1):10–33, April 2007. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2008:NFF

[Rei08] John Reid. The new features of Fortran 2008. ACM Fortran Forum, 27(2):8–21, August 2008. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2010:CNF

[Rei10] John Reid. Coarrays in the next Fortran standard. ACM Fortran Forum, 29(2):10–27, August 2010. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2014:NFF

[Rei14] John Reid. The new features of Fortran 2008. ACM Fortran Forum, 33(2):21–37, August 2014.

CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2017:NFF

[Rei17] John Reid. The new features of Fortran 2015. ACM Fortran Forum, 36(2):3–28, August 2017. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:2018:NFF

[Rei18] John Reid. The new features of Fortran 2018. ACM Fortran Forum, 37(1):5–43, April 2018. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Richardson:2000:HUM

[Ric00] Harvey Richardson. How to use modules. ACM Fortran Forum, 19(1):23–27, April 2000. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ritz:1987:C

[Rit87] Chuck Ritz. Correspondence. ACM Fortran Forum, 6(3):9, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Rolison:1995:DPT

[Rol95] Larry Rolison. Descriptors for pointer target arrays. ACM Fortran Forum, 14(1):39–41, March 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Ross:1995:PSG

[Ros95] Mike Ross. Phrase structure grammar for Fortran. ACM Fortran Forum, 14(3):20, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Runyan:1995:SEC

[Run95] Bob Runyan. Standard Elf “Conservative Fortran” subset? ACM Fortran Forum, 14(3):19, September 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1999:AFU

[RW99a] N. Reid and J. P. Wray. An analysis of Fortran utilization. ACM Fortran Forum, 18(3):2–5, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1999:PSF

[RW99b] N. Reid and J. P. Wray. A prescriptive semantics of Fortran 95. ACM Fortran Forum, 18 (2):2–3, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Sobel:1987:C

[SBDB87] Bernard A. Sobel, Mark D.

Bej, Carrington B. Dixon, and Keith H. Bierman. Correspondence. ACM Fortran Forum, 6 (1):18–27, April 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1991:FEF

[Sch91] Lawrie Schonfelder. The future evolution of Fortran. ACM Fortran Forum, 10(2):16–17, July 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1992:VLS

[Sch92] J. L. Schonfelder. Varying length strings in Fortran. ACM Fortran Forum, 11(2):24–32, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1993:FAO

[Sch93] Lawrie Schonfelder. Fortran 90 is almost object-oriented: some thoughts about Fortran and OOP. ACM Fortran Forum, 12(2):21–22, June 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schulz:1998:OOP

[Sch98] Werner W. Schulz. Objectoriented programming. ACM Fortran Forum, 17(1):11–19, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schwarz:1999:CUM

[Sch99] Michael Schwarz. Compilation using modules with F90. ACM Fortran Forum, 18(3):13–14, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:2001:VPA

[Sch01] J. L. Schonfelder. Variable precision arithmetic: a Fortran 95 module. ACM Fortran Forum, 20(3):2–11, December 2001. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Stern:2002:CCR

[SG02] Ted Stern and Daniel Grimwood. Cascade compilation revisited. ACM Fortran Forum, 21 (1):12–24, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schnebly:1989:Cb

[SH89] Dexter A. Schnebly and Sándor V.

Halász. Correspondence. ACM Fortran Forum, 8(3):11–12, August 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Shenkin:1999:WCF

[She99] Peter S. Shenkin. Writing C functions callable by Fortran and vice versa. ACM Fortran Forum, 18(3):6–12, December 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Shenkin:2000:WOB

[She00] Peter S. Shenkin. Writing objectbased libraries with Fortran and C APLs. ACM Fortran Forum, 19(1):5–12, April 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schlaifer:1983:C

[SHK83] R. Stephen Schlaifer, Albert Hybl, and Steven Kuturoff. Correspondence. ACM ForTec Forum, 2(4):4–7, December 1983. CODEN ???? ISSN 0735-3731.

Siehl:2016:MCF

[Sie16] Michael Siehl. MPMD with Coarray Fortran (2008): an Example Program. ACM Fortran Forum, 35(2):4–15, August 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Singer:1999:FPM

[Sin99]  Jürgen K. Singer.  Is For-

tran portability a myth? crossplatform software development with Fortran. ACM Fortran Forum, 18(1):2–4, April 1999. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schonfelder:1987:PFS

[SM87] Lawrie Schonfelder and Steve Morgan. Pointer functionality sans data-type. ACM Fortran Forum, 6(3):12–26, December 1987. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Shterenlikht:2015:FC

[SMCH15] Anton Shterenlikht, Lee Margetts, Luis Cebamanos, and David Henty. Fortran 2008 coarrays. ACM Fortran Forum, 34 (1):10–30, April 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Smith:1985:SPF

[Smi85] Brian T. Smith. Software parts in Fortran 8X. ACM Fortran Forum, 4(1):33–38, March 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Stone:1998:SSF

[SNB^{+}98] Anthony Stone, Viggo Norum, Harry Bingham, William F. Mitchell, Neil N. Carlson, Douglas Johnson, and Richard Maine. Software in the spotlight: Fortran 90 for Linux. ACM Fortran Forum, 17(1):4, April 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:1997:PMT

[Sny97] Van Snyder. Poor man’s templates in Fortran 90. ACM Fortran Forum, 16(3):11, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:2000:CMF

[Sny00] Van Snyder. Constructing and maintaining Fortran 95 programs using make. ACM Fortran Forum, 19(3):2–8, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:2001:CUI

[Sny01] Van Snyder. Constructive uses for INCLUDE. ACM Fortran Forum, 20(2):2–4, August 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:2004:RFE

[Sny04] Van Snyder. Review of “Fortran 95/2003 Explained, by Michael Metcalf, John Reid, and Malcolm Cohen” OUP, Oxford and New York, 2004. ACM Fortran Forum, 23(2):12, August 2004. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:2015:EHF

[Sny15] Van Snyder. Exception handling in Fortran. ACM Fortran Forum, 34(1):5–9, April 2015. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Snyder:2016:EHF

[Sny16] Van Snyder. Exception handling in Fortran. ACM Fortran Forum, 35(2):24–27, August 2016. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Soffa:1999:ALM

[Sof99] Mary Lou Soffa. Award for Loren Meissner. ACM Fortran Forum, 18(1):1, April 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Stiller:2000:MM

[SSH^{+}00] Joerg Stiller, Ted Stern, Phillip Helbig, Neil Carlson, Glenn Carver, David Vowles, and Viktor Decyk. More on modules. ACM Fortran Forum, 19(2):24– 28, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Stewart:2003:MLD

[Ste03] G. W. Stewart. Memory leaks in derived types revisited. ACM Fortran Forum, 22(3):25–27, December 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Simpson:1993:RMF

[SVM^{+}93] Dan Simpson, Sam Valentine, Richard Mitchell, Lulu Liu, and Rod Ellis. Recoup — maintaining Fortran. ACM Fortran Forum, 12(3):26–32, September 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Schnebly:1989:Ca

[SWM^{+}89] Dexter A. Schnebly, Bruce W.

Walker, Richard R. Moore, John R. Rice, J. H. Matheny, and George Shaw. Correspondence. ACM Fortran Forum, 8(1):21– 32, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Scott:2001:MMF

[SWM01] Gary Scott, Tony T. Warnock, and Patrick F. McGehearty. Maintenance mode for the Fortran Standard? ACM Fortran Forum, 20(1):25–26, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Taylor:1997:SQS

[Tay97] Howard L. Taylor. Some questions and suggestions concerning Fortran 90/95. ACM Fortran Forum, 16(1):19–20, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

PCF:1991:SIP

[The91] The Parallel Computing Forum.

Special issue: PCF parallel Fortran extensions. ACM Fortran Forum, 10(3):1–57, September 1991. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Tholen:2000:FH

[TLBC00] Dave Tholen, Steve Lionel, Jos Bergervoet, and Ewan Cunningham. Fortran in the heavens.

ACM Fortran Forum, 19(3):12, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). Notes that an asteroid has been named “Fortran”.

Tinetti:2012:FLS

[TM12] Fernando G. Tinetti and Mariano Méndez. Fortran legacy software: source code update and possible parallelisation issues. ACM Fortran Forum, 31 (1):5–22, April 2012. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Thornburg:1988:C

[TSSL88] Jonathan Thornburg, Houston H. Stokes, Nora H. Sabelli, and Robert Loveless. Correspondence. ACM Fortran Forum, 7 (1):9–16, April 1988. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Tanner:1983:C

[TV83] Dan Tanner and David M. Vallance. Correspondence. ACM ForTec Forum, 2(3):5–8, September 1983. CODEN ???? ISSN 0735-3731.

Vallance:1983:FCS

[Val83] D. M. Vallance. Fortran 77 compiler survey. ACM ForTec Forum, 2(1):9–12, March 1983. CO-DEN ???? ISSN 0735-3731.

VanTuyl:1984:EF

[Van84] Robert R. Van Tuyl. On evolution of Fortran. ACM Fortran Forum, 3(3):39–42, December 1984. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Vaught:2000:GFF

[Vau00] Andy Vaught. Gnu Fortran — it’s free crunch time. ACM Fortran Forum, 19(3):13–14, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Verma:1990:TMI

[Ver90] Ajit Kumar Verma. Towards more informative Fortran compilers. ACM Fortran Forum, 9 (1):29–30, March 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Vowels:1999:WPF

[Vow99] R. A. Vowels. Writing portable Fortran procedures. ACM Fortran Forum, 18(2):8–11, August 1999. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Vowels:2001:TGC

[Vow01] R. A. Vowels. A truly generic CMPLX function. ACM Fortran Forum, 20(3):18–19, December 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Vowels:2002:RVS

[Vow02] R. A. Vowels. Returning a varying string. ACM Fortran Forum, 21(1):8–11, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1982:CLI

[Wag82a] Jerrold L. Wagener. Chairman’s letter: Introducing ForTec Forum. ACM ForTec Forum, 1(1): 1, July 1982. CODEN ???? ISSN 0735-3731.

Wagener:1982:E

[Wag82b] Jerrold L. Wagener. From the Editor: Volunteers needed; European Subcommittee proposed; quotes. ACM ForTec Forum, 1 (2):1, December 1982. CODEN ???? ISSN 0735-3731.

Wagener:1983:SIS

[Wag83a] Jerrold L. Wagener. Short INTE-GERS in standard Fortran 77. ACM ForTec Forum, 2(3):9–10, September 1983. CODEN ???? ISSN 0735-3731.

Wagener:1983:EFP

[Wag83b] Jerrome L. Wagener. From the Editor: Fortran 77 on personal computers?; Fortran 77 compiler survey; Fortran at NCC. ACM ForTec Forum, 2(2):1, June 1983. CODEN ???? ISSN 0735-3731.

Wagener:1984:SWT

[Wag84] Jerrold L. Wagener. Status of work toward revision of programming language Fortran. ACM ForTec Forum, 3(2):1–42, June 1984. CODEN ???? ISSN 0735-3731. Co-published as ACM SIGNUM Newsletter, v. 19, no. 3 (July 1984).

Wagener:1985:IFG

[Wag85a] J. L. Wagener. Is Fortran getting too big? ACM Fortran Forum, 4(2):16, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1985:FGT

[Wag85b] Jerrold L. Wagener. Is Fortran getting too big? ACM Fortran Forum, 4(2):16, June 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1993:SIF

[Wag93]  Jerrold L. Wagener.  Spe-

cial issue: Fortran 90 errata, amendments and interpretations: Progress to date. ACM Fortran Forum, 12(1):1–48, March 1993. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1995:FFI

[Wag95] Jerrold Wagener. Fortran 95 and Fortran 2000: International working group WG5, US Fortran committee X3J3, Fortran interoperability with C, pure and elemental procedures in Fortran 95, dummy arguments with target attribute. ACM Fortran Forum, 14(4):7–12, December 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1997:FRa

[Wag97a] Jerry Wagener. Fortran reflections. ACM Fortran Forum, 16 (1):1, April 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1997:FRb

[Wag97b] Jerry Wagener. Fortran reflections. ACM Fortran Forum, 16 (2):1, August 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:1997:FRc

[Wag97c] Jerry Wagener. Fortran reflections. ACM Fortran Forum, 16 (3):1, December 1997. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagener:2003:FPA

[Wag03] Jerry Wagener. Fortran photo archive. ACM Fortran Forum, 22(2):18, August 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Walster:2000:UII

[Wal00] G. William Walster. The use and implementation of interval data types in Fortran. ACM Fortran Forum, 19(2):2–15, August 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). Describes the interval arithmetic support in the Sun Microsystems Forte Developer 6 Fortran 95 (f95) compiler.

Walter:2003:GNV

[WCH^{+}03] Wolfgang Walter, Norman S. Clerman, Dick Hendrickson, Richard Maine, Van Snyder, John Reid, and David Muxworthy. Germany’s negative vote. ACM Fortran Forum, 22(1):11– 21, April 2003. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Whitlock:2000:WDK

[Whi00] Stan Whitlock. When default KIND isn’t good enough. ACM Fortran Forum, 19(3):15–16, December 2000. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Wada:1983:C

[WM83] Hideo Wada and Austin J. Maher. Correspondence. ACM ForTec Forum, 2(1):4–6, March 1983. CODEN ???? ISSN 0735-3731.

Weaver:1992:CCP

[WRV^{+}92] Dick Weaver, John Reid, Daniel J.

Vasicek, Lawrie Schonfelder, and Keith Bierman. Correspondence: Concerning PCF report in Fortran Forum. ACM Fortran Forum, 11(1):12–16, March 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).